### v1

# PADetBench: Towards Benchmarking Physical Attacks against Object Detection

Abstract
Physical attacks against object detection have gained increasing attention due to their significant practical implications. However, conducting physical experiments is extremely time-consuming and labor-intensive. Moreover, physical dynamics and cross-domain transformation are challenging to strictly regulate in the real world, leading to unaligned evaluation and comparison, severely hindering the development of physically robust models. To accommodate these challenges, we explore utilizing realistic simulation to thoroughly and rigorously benchmark physical attacks with fairness under controlled physical dynamics and cross-domain transformation. This resolves the problem of capturing identical adversarial images that cannot be achieved in the real world. Our benchmark includes 20 physical attack methods, 48 object detectors, comprehensive physical dynamics, and evaluation metrics. We also provide end-to-end pipelines for dataset generation, detection, evaluation, and further analysis. In addition, we perform 8064 groups of evaluation based on our benchmark, which includes both overall evaluation and further detailed ablation studies for controlled physical dynamics. Through these experiments, we provide in-depth analyses of physical attack performance and physical adversarial robustness, draw valuable observations, and discuss potential directions for future research. Codebase: GitHub Repository.
1 Introduction
Deep neural networks (DNNs) have achieved remarkable success in various research fields and industrial applications, such as computer vision [65], natural language processing [64], and speech recognition [61]. However, previous studies [78, 27, 2, 41, 3] have shown that DNNs are susceptible and vulnerable to adversarial attacks, which can be divided into digital attacks and physical attacks according to different attack domains in computer vision tasks. Technically, digital attacks are usually performed by adding imperceptible perturbations to the input image of victim models after imaging, while physical attacks can cheat DNNs by directly modifying the physical properties of the targets before imaging, such as changing the texture of the object [77, 114], and adding stickers [96, 44]. Intuitively, physical attacks are more dangerous and practical than digital attacks because they can be easily implemented in the real world with fewer prerequisites, which arouses significant concerns, especially in safety-critical applications like autonomous driving [92, 6], security surveillance [63, 95], remote sensing [94, 48], and so on.
Object detection is a fundamental and pragmatic task in computer vision, which has been widely deployed in various intelligent systems [126, 113]. Consequently, most of the physical attacks aim to fool object detectors in real-world scenarios, and the physical adversarial robustness of object detection models has attracted increasing attention in recent years. However, the absence of regulated and easy-to-follow benchmarks hinders the development of physical attack and physically robust detection methods. The main reasons for discouraging the development of physical attack benchmarks are as follows. 1) Time-consuming and expensive: To evaluate the physical attack performance and physical adversarial robustness of object detection models, researchers need to conduct a large number of experiments in the real world, which is time-consuming and expensive. 2) Physical dynamics align: To ensure comparison fairness, physical dynamics should be strictly controlled and kept the same, while this can’t be achieved in real-world scenarios since it is impossible to take two identical pictures. 3) Cross-domain loss: Physical attacks are commonly conducted by elaborating conspicuous adversarial patterns that should be able to survive after transforming from the physical to the digital domain, while this cross-domain loss is uncontrollable. 4) Hard to compare: With the progressive evolution of physical attacks, physical adversarial perturbations have evolved from 2D space to 3D space, which builds a barrier to fairly compare different kinds of physical attack methods. As the above challenges exist, it is tough to effectively verify the physical attack efficacy and physical adversarial robustness of object detection models without thorough evaluation and impartial comparisons. As a result, the researchers can not identify the actual progress of physical adversarial attacks and physical adversarial robustness development, which may gear down the progress of the research field.
In this paper, we propose a fair and rigorous benchmarking of physical attacks on object detection models. We utilize CARLA [17], a highly realistic autonomous driving simulator, to benchmark physical attacks under consistent physical dynamics such as weather, viewing angle, and distance. These conditions are challenging to align for impartial comparisons in the real world due to the impossibility of capturing two identical photographs. Our benchmark includes 20 physical attack methods, clean and random noised objects for comparison, 48 object detectors, diverse physical dynamics, evaluation metrics from different perspectives, and a comprehensive pipeline for evaluating physical attacks and subsequent analysis. We conduct exhaustive evaluation of each physical attack against a wide range of object detectors under various physical dynamic conditions, using multiple evaluation metrics. Following our experimental procedures, we delve into detailed analyses concerning the performance of physical attacks and the robustness against physical adversarial actions. We highlight intriguing findings and discuss potential avenues for future research.
In summary, our contributions are as follows:
We propose a robust and equitable benchmark for physical attacks against object detection models. This benchmark deeply explores the potential of real-world simulators to consistently evaluate physical attacks under a variety of continuous physical dynamics.
We offer 8064 groups of evaluation on 20 physical attacks against 48 object detectors, which are conducted under continuous physical dynamics with diverse evaluation metrics, providing a holistic perspective on the performance of attack and detection.
We delve into detailed analyses of the fooling efficacy of physical attacks and the robustness of object detection models against these attacks. Through these analyses, we highlight valuable findings and discuss potential avenues for future research in this area.
2 Related work
2.1 Object detection
Object detection serves as a cornerstone in computer vision, with the primary objective of identifying and localizing objects within images or video sequences. With the advent of deep learning, DNNs have been widely adopted for object detection, significantly improving performance. R-CNN [26] and its successors [25, 70, 56, 66, 99, 106, 75], have made substantial strides in improving detection speed and accuracy by introducing region proposal networks and shared convolution computations. Single Shot MultiBox Detector (SSD) [55] and You Only Look Once (YOLO) [67] series [67, 68, 69, 1, 38, 43, 86, 37, 87, 85] further accelerate the detection process by eliminating the need for region proposals and performing detection in a single pass, making them suitable for real-time applications. Recently, more advanced architectures based on transformers have been proposed, such as DETR [7], DAB-DETR [52], ViTDet [47], DINO [104], and Co-DETR [125], which further push the boundaries of object detection performance by leveraging attention mechanisms. Despite these advancements, object detection in adversarial environments remains a challenging task, necessitating continuous research and development in this area.
2.2 Physical attack
Kurakin et al. [41] first illustrate that machine learning systems remain susceptible to adversarial examples, even within these physical world contexts. They substantiate this by introducing adversarial images, captured via a cell phone camera, which still can dramatically degrade the performance of intelligent vision systems. Brown et al. [2] first design adversarial perturbations in patch form, i.e., adversarial patch, which localizes adversarial perturbations to a specific region of the image without imperceptibility constraints. Adversarial patches are more practical and effective in the real world, as they can be easily printed and attached to objects to fool detectors [73, 81, 100, 124, 122, 91, 121, 34, 112, 72, 35, 29]. Nonetheless, the adversarial patches mentioned above are overly noticeable, potentially arousing the suspicion of the victims. To mitigate this concern, scholars have suggested the generation of adversarial patches in natural styles to circumvent drawing attention [36, 32, 28]. Beyond adversarial patches, the academic community has also investigated physical perturbations in a multitude of forms, including light [31, 98], viewpoint [15], and 3D objects [51], among others. To reconcile the disparity between 2D perturbations and 3D physical space, the concept of adversarial perturbations has been expanded into the 3D realm [111, 88, 76, 77, 117]. This extension has proven to be more efficacious and applicable in real-world scenarios. The considerable variation in physical perturbations and experimental settings presents a hurdle to the equitable comparison of disparate physical attack methodologies.
2.3 Robustness benchmark
The benchmarking of adversarial attacks is crucial for evaluating the robustness of DNNs-based models and further developing more robust ones. To address the issue of overestimated robustness, [12] first establish a standardized benchmark of adversarial robustness, which accurately reflects the robustness of the considered models within a reasonable computational budget. Wu et al. [97] build a comprehensive benchmark for backdoor attacks in image classification models. [60] provides an easy-to-use benchmark to assess the performance of object detection models under conditions of deteriorating image quality, such as image distortions or diverse weather conditions instead of adversarial perturbations. In contrast, [115] benchmark the adversarial robustness of image classifiers in black-box settings. Except for RGB images, Dong et al. [14] benchmark the robustness of 3D object detection to common corruptions from weather, sensor, motion, etc., in both LiDAR and camera data. Li et al. [45] towards benchmarking and assessing the visual naturalness of physical adversarial perturbations. Hingun et al. [30] construct a large-scale benchmark for evaluating adversarial patches with traffic sign dataset. CARLA [17], as a realistic autonomous driving simulator, has been progressively involved in the research of physical adversarial robustness. [62] presents CARLA-GEAR: a dataset generator for evaluating the adversarial robustness of vision models. The work in [109] put forth a pipeline for instance-level data generation, leveraging the CARLA simulator. Utilizing this pipeline, they created the DCI dataset and carried out experiments involving three detectors and three physical attacks. In summary, there is still a lack of a comprehensive and rigorous benchmark for physical attacks against object detection models. We aim to fill the gap in this work and significantly lower the threshold of this path with easy-to-follow instructions and codebase.
3 PADetBench
The benchmark encompasses four integral facets: datasets, physical attack implementations, object detection models, and comprehensive evaluation & analysis procedures. From a technical standpoint, we have engineered each constituent of the benchmark as modular, end-to-end pipelines within our publicly accessible codebase, ensuring straightforward adoption and replication.
3.1 Dataset Generation
It is common to use COCO [50], PASCAL VOC [19], KITTI [23], etc., as benchmark datasets for object detection. However, these datasets are ill-suited for assessing physical attacks since they are static and lack the flexibility required to create manipulated, real-world adversarial scenarios. Physical attacks typically entail altering the physical attributes of objects before capturing their images. To fairly and accurately evaluate and compare such attacks, experiments necessitate applying perturbations in real-world conditions with controlled physical dynamics, which are excessively time-consuming, labor-intensive, and theoretically infeasible. Simulated environments, like CARLA [17], present a viable solution to these obstacles by enabling the straightforward manipulation of physical dynamics through configurable parameters.
This work contributes an end-to-end pipeline for dataset generation within our codebase, significantly streamlining the dataset generation process and enhancing research productivity. Our pipeline prioritizes user-friendliness, enabling researchers to swiftly generate datasets embodying diverse physical conditions through a concise series of steps. These conditions encompass variations in weather, viewing angles, and distances, along with the capacity to impose physical perturbations on objects. Comprehensively, our pipeline supports over 10 distinct environments ranging from downtowns to small towns and rural landscapes (detailed listing available in the Appendix), coupled with a library of more than 40 vehicles and 40 pedestrian models, all customizable concerning their hues and surface textures. It further integrates continuous manipulation of physical dynamics such as fluctuating weather patterns, precise sun positioning, and flexible camera placements concerning both location and orientation. To ensure accessibility, we accompany the pipeline with step-by-step guidelines for personalizing object perturbations and seamlessly integrating these modifications within CARLA’s [17] simulation framework.
Our benchmark comprises three categories of datasets: a clean dataset serving as a control group, a dataset with random noise perturbations, and several datasets featuring adversarial perturbations generated through various attack methodologies. To ensure fair comparisons, scene compositions and camera perspectives are meticulously synchronized and regulated across all datasets, achievable effortlessly through our provided pipeline.
Moreover, our pipeline facilitates the automatic generation of supplementary annotations, including 2D and 3D bounding boxes, depth maps, and instance segmentation maps. Consequently, our benchmark extends its utility beyond 2D object detection, also catering to tasks like 3D object detection, instance segmentation, depth estimation, and more, thereby enhancing the scope of research and application in computer vision.
3.2 Physical attacks
We select 20 representative physical attack methods, in which FCA [88], DTA [76], ACTIVE [77], 3D2Fool [114], POOPatch [11], RPAU [54], APPA [48], and CAMOU [111], are for vehicle detection, while DAP [29], AdvPattern [95], UPC [36], NatPatch [32], MTD [13], AdvCaT [33], AdvTexture [34], AdvTshirt[101], AdvPatch [81], LAP [79], InvisCloak [100], and AdvCam [18], are for person detection. The corresponding physical perturbations of these methods are imported into Unreal Engine 4 for CARLA (CarlaUE4) [17], as shown in the physical attacks part of Fig. 1, to generate the physical adversarial datasets. We adhere to two principles similar to [97] when selecting physical attacks. First, the methods are representative or advanced in the research field, which can serve as baseline and state-of-the-art (SOTA) methods for comparison, respectively. Second, physical attacks are easily conducted and with reproducible performance, which can be conveniently followed and reproduced by other researchers. Since our benchmark evaluates physical attacks based on their crafted perturbations, novel physical attack methods can be easily integrated into the benchmark by following the provided pipeline. We will continue to update the physical attacks in the benchmark to keep pace with the latest research progress.
Backbone | Category | Detectors |
CNN | One-stage | ATSS[107], AutoAssign[119], CenterNet[118], CentripetalNet[16] |
CornerNet[42], DDOD[10], DyHead[99], EfficientNet[80], FCOS[82], | ||
FoveaBox[40], FreeAnchor[110], FSAF[120], GFL[46],LD[116], NAS-FPN[24], | ||
PAA[39], RetinaNet[49], RTMDet[58], TOOD[20], VarifocalNet[105], | ||
YOLOX[22], YOLOv5[38], YOLOv6[43], YOLOv7[86], YOLOv8[37] | ||
Two-stage | Faster R-CNN[70], Cascade R-CNN[5], Cascade RPN[84], | |
Double Heads[99], FPG[8],Grid R-CNN[56], Guided Anchoring[89], | ||
HRNet[74], Libra R-CNN[66], PAFPN[53],RepPoints[102], | ||
Res2Net[21], ResNeSt[103], SABL[90], Sparse R-CNN[75] | ||
Transformer | - | DETR[7], Conditional DETR[59], DDQ[108], DAB-DETR[52], |
Deformable DETR[123], DINO[104],PVT[93], PVTv2[93] |
3.3 Object detectors
We choose 48 object detectors in the same principles as choosing physical attack methods, covering mainstream object detectors, such as YOLO series [38, 43, 86, 37, 22] (One-stage) and R-CNN series [26, 25, 70, 4, 75], which are based on CNN. Except for canonical detectors, we also include transformer-based detectors, such as DETR [7], Conditional DETR [59], Deformable DETR [123], DAB-DETR [52], and DINO [104]. All the selected detectors are listed in Table 1 according to their characteristics. Our benchmark provides the end-to-end pipeline for object detection evaluation based on MMDetection [9]. Consequently, it is convenient to integrate new detectors into the benchmark, and the benchmark can also be easily extended to evaluate other vision tasks, such as 3D object detection, instance segmentation, and depth estimation.
3.4 Evaluation and analysis
To rigorously assess the efficacy of physical attacks on object detection systems, we furnish baseline datasets: clean datasets (without perturbations) and those infused with randomized noise (incorporating arbitrary disturbances in -bounded space). This dual-baseline approach sets the stage for a thorough examination.
Quantifying performance entails employing evaluation metrics that consider both detection accuracy and the performance of the attacks. These metrics comprise by several widely adopted indicators, including average recall (AR) and Attack Success Rate (ASR). AR is calculated as the average of the recall rates:
where TP and FN denote the true positive and false negative counts of the detector, respectively. AR signifies the proportion of correctly identified instances relative to the actual total in the dataset, offering an intuitive gauge of how physical attacks degrade the detection capability of a given target. On the other hand, ASR quantifies the effectiveness of the adversarial perturbations, calculated as:
where and denote the average recall rates of the attack and clean datasets, respectively. ASR provides a direct measure of the extent to which the attacks undermine the detector’s confidence in its predictions.
Furthermore, we enhance our codebase by incorporating several ready-to-use explainability visualization tools, facilitating deeper insights into model behavior. These include Grad-CAM [71] for visualizing the regions of input data that contribute most to the model’s prediction, Shapley value [57] to quantify the individual feature contributions, and t-SNE [83] for reducing dimensionality and visualizing high-dimensional data in a more interpretable manner. These additions empower users to conduct comprehensive analyses beyond mere performance evaluation.
4 Experiments
4.1 Experimental setup
Datasets. 1) Overall experiments. We generate overall datasets with 3 objects, 10 weather conditions, 2 altitude angles, 8 azimuth angles, 5 radius values, 3 spawn points, and 20 physical perturbations, i.e., 7200 samples () for each attack method, in which the physical dynamics are strictly aligned and controlled for impartial comparison. 2) Ablation Studies. We conduct in-depth examinations to explore the individual impact of core physical dynamics: weather conditions, venue, camera distance, azimuth angle, altitude angle within a hemispherical space. Accomplishing this involves generating focused sub-benchmarks, each consisting of 100 samples.
Physical attacks. we generate 24 datasets, including 20 physically noised datasets that correspond to 20 physical attacks, an extra 2 clean datasets and 2 randomly noised datasets for comparison of vehicle detection and person detection, respectively.
Object detectors. We evaluate 48 object detectors, as shown in Table 1, by integrating MMDetection [9] into our evaluation pipeline, in which the selected detectors are trained on the COCO [50] dataset.
Therefore, we conduct a total of 8064 () groups of the experiment, which are conducted with NVIDIA Geforce 4090.
4.2 Overall experiments and analysis
We present the comprehensive results of vehicle detection against physical attacks in Fig. 2. Additionally, Fig. 3 and Fig. 4 show visualized analyses of the experimental results from detection and attack perspectives, respectively. Experimental results of person detection and corresponding detailed numerical results are listed in the Appendix. From these analyses, several key observations emerge:
Detection perspective: 1) Vehicle detection performance is significantly impacted by physical attacks, with the average recall rates of detectors decreasing up to 50%, as shown in Fig. 4. However, pedestrian detection performance is less affected regarding various attacks, with the average recall rates of detectors decreasing by less than 20%, as shown in Fig. 3. The potential reason is that the stronger physical perturbations are optimized with consideration of 3D space and accommodate more complex physical dynamics, while physical attacks aiming to fool person detectors are commonly performed with optimized 2D patches, which work well in particular physical dynamics, as shown in the ablation experiments, which empirically demonstrate the pressing need and necessity of a comprehensive and rigorous benchmark for physical attacks. 2) For different detectors, the performance varies significantly, with some detectors exhibiting superior robustness against physical attacks like EfficientNet, YOLO series and RTMDet of one-stage detectors. Besides, DDQ also shows significant adversarial robustness among transformer-based detectors. While the rest of the detectors show varying lower levels of robustness, the SOTA detection performance is not necessarily indicative of adversarial robustness. Consequently, the benchmark also serves as a robustness indicator.
Attack perspective: 1) For vehicle detection, different physical attacks exhibit varying levels of effectiveness, with some attacks achieving ASR values exceeding 70% like ACTIVE, and others failing to surpass 20%. Most of the physical attacks hard to fool the latest SOTA detectors, such as EfficientNet, YOLO series, and RTMDet. This phenomenon is caused by the victim models of the attack method lagging behind the development of the detection method, which also motivates us to fill this gap. 2) For person detection, the ASR values of physical attacks are generally lower than those for vehicle detection, with the majority of attacks achieving ASR values below 20%, even lower than 0% (see more explanation in the Appendix). The relatively strongest attack method is AdvTexture, which elaborates on a 2D patch but with tricks for 3D space. This also demonstrates the gap between 2D perturbations and 3D physical space.
4.3 Ablation experiments and analysis
Except for the overall experiments, we also conduct ablation experiments to investigate the impact of physical world factors. We show the results of 3 physical dynamics, including weather, distance, and camera viewing angle, in Fig. 5 and Fig. 6, respectively. More experiments on other dynamics are provided in the Appendix. From these analyses, several key observations emerge: 1) Physical attack performance can be easily swayed by physical dynamics. This phenomenon is consistent with our intuition and emphasizes the importance of strictly aligning physical dynamics when evaluating physical attacks, which are often underestimated by previous works. 2) We also observe a gap between the attack performance of our benchmark and the reported performance in the original papers (supplemented in Appendix). Two reasons may contribute to this gap: the first is the adopted SOTA detectors in our benchmark, which are more robust than the victim models in the original papers, and the second is that our benchmark provides more comprehensive and strict evaluation datasets and physical dynamics, which are more challenging for the attack methods. Please refer to the Appendix for more experiments, detailed analysis and discussion.
5 Discussion
Limitations. For now, PADetBench mainly focuses on the physical attacks on object detection models. In the future, we will extend the benchmark to evaluate other vision tasks, such as instance segmentation, 3D object detection, and depth estimation.
Potential impacts. 1) Positive impacts: In-depth understanding of the robustness of object detection models against physical attacks, which can guide the development of more robust models. 2) Negative impacts: The benchmark may be misused to conduct physical attacks in real life, which may pose a threat to the security-critical applications involving intelligent visual perception.
6 Conclusion
In conclusion, this work aims to meet the critical need for a standardized and rigorous benchmark for physical attacks on object detection models. Our benchmark leverages the CARLA simulator for consistent physical dynamics and includes 20 attack methods, 48 detectors, comprehensive dynamics, and metrics from different perspectives. It also provides end-to-end pipelines for dataset generation, attack and detection evaluation, and experimental results analysis. Our extensive evaluation and analyses provide valuable insights for future research directions. The supporting codebase is publicly available and we will continuously supplement it, aiding the development of physically robust object detection methods.
7 Acknowledgment
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes]
Did you discuss any potential negative societal impacts of your work? [Yes]
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Please refer to the GitHub repository: https://github.com/JiaweiLian/Benchmarking_Physical_Attack for the code and instructions.
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A] This benchmark does not involve a training process.
Did you report error bars (e.g., to the random seed after running experiments multiple times)? [N/A]
Did you include the total amount of computing and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
- (a)
Did you mention the license of the assets? [Yes] Please refer to the Appendix for the license.
Did you include any new assets either in the supplemental material or as a URL? [Yes] Please refer to the GitHub repository (https://github.com/JiaweiLian/Benchmarking_Physical_Attack) for details.
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Additional content of the Benchmark
A.1 Mini-test
We kindly invite the reviewers and readers to participate in a mini-test to discriminate the real-world images and the simulated images as shown in Fig. 7, the answer is revealed in its caption.
Town | Description |
Town1 | A small, simple town with a river and several bridges. |
Town2 | A small simple town with a mixture of residential and commercial buildings. |
Town3 | A larger, urban map with a roundabout and large junctions. |
Town4 | A small town embedded in the mountains with a special "figure of 8" infinite highway. |
Town5 | Squared-grid town with cross junctions and a bridge. |
It has multiple lanes per direction. Useful to perform lane changes. | |
Town6 | Long many lane highways with many highway entrances and exits. It also has a Michigan left. |
Town7 | A rural environment with narrow roads, corn, barns and hardly any traffic lights. |
Town8 | Secret "unseen" town used for the Leaderboard challenge. |
Town9 | Secret "unseen" town used for the Leaderboard challenge. |
Town10 | A downtown urban environment with skyscrapers, residential buildings and an ocean promenade. |
Town11 | A Large Map that is undecorated. |
Serves as a proof of concept for the Large Maps feature. | |
Town12 | A Large Map with numerous different regions, |
including high-rise, residential and rural environments. |
Full list of the optional maps, where Town8 and Town9 are unseen for competition. Please refer to CARLA [17] documentary for more details.
A.2 Corresponding config files of the selected detectors
The corresponding config files of the selected detectors are listed in Table 3.
Number | Config Files | Detectors |
1 | atss_r50_fpn_1x_coco | ATSS[107] |
2 | autoassign_r50-caffe_fpn_1x_coco | AutoAssign[119] |
3 | centernet-update_r50-caffe_fpn_ms-1x_coco | CenterNet[118] |
4 | centripetalnet_hourglass104_16xb6-crop511-210e-mstest_coco | CentripetalNet[16] |
5 | cornernet_hourglass104_10xb5-crop511-210e-mstest_coco | CornerNet[42] |
6 | ddod_r50_fpn_1x_coco | DDOD[10] |
7 | atss_r50_fpn_dyhead_1x_coco | DyHead[99] |
8 | retinanet_effb3_fpn_8xb4-crop896-1x_coco | EfficientNet[80] |
9 | fcos_x101-64x4d_fpn_gn-head_ms-640-800-2x_coco | FCOS[82] |
10 | fovea_r50_fpn_4xb4-1x_coco | FoveaBox[40] |
11 | freeanchor_r50_fpn_1x_coco | FreeAnchor[110] |
12 | fsaf_r50_fpn_1x_coco | FSAF[120] |
13 | gfl_r50_fpn_1x_coco | GFL[46] |
14 | ld_r50-gflv1-r101_fpn_1x_coco | LD[116] |
15 | retinanet_r50_nasfpn_crop640-50e_coco | NAS-FPN[24] |
16 | paa_r50_fpn_1x_coco | PAA[39] |
17 | retinanet_r50_fpn_1x_coco | RetinaNet[49] |
18 | rtmdet_s_8xb32-300e_coco | RTMDet[58] |
19 | tood_r50_fpn_1x_coco | TOOD[20] |
20 | vfnet_r50_fpn_1x_coco | VarifocalNet[105] |
21 | yolov5_l-p6-v62_syncbn_fast_8xb16-300e_coco | YOLOv5[38] |
22 | yolov6_l_syncbn_fast_8xb32-300e_coco | YOLOv6[43] |
23 | yolov7_l_syncbn_fast_8x16b-300e_coco | YOLOv7[86] |
24 | yolov8_l_syncbn_fast_8xb16-500e_coco | YOLOv8[37] |
25 | yolox_l_fast_8xb8-300e_coco | YOLOX[22] |
26 | faster-rcnn_r50_fpn_1x_coco | Faster R-CNN[70] |
27 | cascade-rcnn_r50_fpn_1x_coco | Cascade R-CNN[5] |
28 | cascade-rpn_faster-rcnn_r50-caffe_fpn_1x_coco | Cascade RPN[84] |
29 | dh-faster-rcnn_r50_fpn_1x_coco | Double Heads[99] |
30 | faster-rcnn_r50_fpg_crop640-50e_coco | FPG[8] |
31 | grid-rcnn_r50_fpn_gn-head_2x_coco | Grid R-CNN[56] |
32 | ga-faster-rcnn_x101-32x4d_fpn_1x_coco | Guided Anchoring[89] |
33 | faster-rcnn_hrnetv2p-w18-1x_coco | HRNet[74] |
34 | libra-retinanet_r50_fpn_1x_coco | Libra R-CNN[66] |
35 | faster-rcnn_r50_pafpn_1x_coco | PAFPN[53] |
36 | reppoints-moment_r50_fpn_1x_coco | RepPoints[102] |
37 | faster-rcnn_res2net-101_fpn_2x_coco | Res2Net[21] |
38 | faster-rcnn_s50_fpn_syncbn-backbone+head_ms-range-1x_coco | ResNeSt[103] |
39 | sabl-faster-rcnn_r50_fpn_1x_coco | SABL[90] |
40 | sparse-rcnn_r50_fpn_1x_coco | Sparse R-CNN[75] |
41 | detr_r50_8xb2-150e_coco | DETR[7] |
42 | conditional-detr_r50_8xb2-50e_coco | Conditional DETR[59] |
43 | ddq-detr-4scale_r50_8xb2-12e_coco | DDQ[108] |
44 | dab-detr_r50_8xb2-50e_coco | DAB-DETR[52] |
45 | deformable-detr_r50_16xb2-50e_coco | Deformable DETR[123] |
46 | dino-4scale_r50_8xb2-12e_coco | DINO[104] |
47 | retinanet_pvt-t_fpn_1x_coco | PVT[93] |
48 | retinanet_pvtv2-b0_fpn_1x_coco | PVTv2[93] |
Appendix B Additional content of the Experiments
B.1 Genetated data for the ablation studies
We provide the generated data samples for the ablation studies in Fig. 8.
B.2 Detailed illustration of the gap between the attack performance of our benchmark and the reported performance in the original papers
Our benchmark addresses two important aspects that were not considered in the original papers:
1.
Strict physical dynamics: Our benchmark leverages the CARLA simulator, which provides a realistic and consistent physical environment for evaluating physical attacks on object detection models. This ensures that the attacks are evaluated under conditions that closely resemble the physical world, allowing for more accurate assessment of the robustness of the models.
2.
Evaluation in 3D space perspective: Unlike the original papers, which mainly focus on 2D evaluation, our benchmark evaluates the attack performance in a 3D space perspective. This is crucial because physical attacks can exploit vulnerabilities in the 3D structure of the objects, such as comprehensive and continuous view, scale, distance, etc. By considering the 3D space perspective, our benchmark provides a more comprehensive evaluation of the models’ robustness against physical attacks.
By addressing these two aspects, our benchmark provides a more realistic and comprehensive evaluation of the attack performance of object detection models, bridging the gap between the reported performance in the original papers and the real-world scenarios.
B.3 Supplemented experiments analysis and discussion
B.3.1 Detection Perspective
Vehicle Detection Perspective:
Physical attacks on vehicle detection systems pose a substantial challenge due to the specialized nature of the perturbations crafted to deceive these models. These attacks can lead to a drastic decline in average recall rates, reaching as low as 50%. This high level of vulnerability is largely attributed to the complex dynamics in the 3D environment where vehicles operate. Physical attacks on vehicle detectors exploit this three-dimensional context, introducing perturbations that consider real-world factors such as lighting, perspective, occlusion, and motion, making them more effective in disrupting the model’s performance.
On the other hand, pedestrians, operating in a somewhat simpler 2D plane, seem to be less affected by similar adversarial attacks, with a decrease in average recall rates of less than 20%. Adversarial examples targeting pedestrian detection typically involve 2D patches, which might be more straightforward to apply in specific scenarios but may not account for the full range of real-world complexities. As a result, there is an urgent demand to establish a comprehensive and stringent benchmark to systematically evaluate the resilience of these models against physical attacks, facilitating research and development towards more secure systems.
Pedestrian Detection Perspective:
In contrast to vehicle detection, pedestrian detectors exhibit a certain level of inherent robustness, potentially due to the simpler constraints imposed on the recognition process. Nevertheless, as seen across various detectors, the extent of this robustness varies widely. Models like EfficientNet, YOLO series, RTMDet (one-stage detectors), and DDQ (transformer-based detectors) demonstrate commendable resistance to physical attacks. The superior performance of DDQ could be linked to the attention mechanisms inherent to transformer architectures, which are capable of capturing global spatial dependencies, thus mitigating the impact of adversarial perturbations.
However, it is evident from the benchmark results that not all state-of-the-art (SOTA) detectors offer comparable adversarial robustness. Many detectors exhibit varying degrees of vulnerability, indicating that peak accuracy in standard detection tasks does not automatically guarantee resilience against adversarial threats. Consequently, this benchmarking framework not only identifies areas of weakness for refinement but also contributes to a better understanding of the interplay between detection performance and adversarial robustness in real-world deployments.
In conclusion, understanding and mitigating the effects of physical attacks in both vehicle and pedestrian detection domains can greatly benefit deep learning and computer vision research. By developing more robust models resistant to such attacks, we can enhance the safety and reliability of autonomous systems that rely on accurate object detection, ultimately fostering advancements in the fields of automotive technology, smart city infrastructure, and robotics. Furthermore, this benchmark would encourage researchers to explore defensive techniques and novel architectures that better withstand both digital and physical adversarial threats, pushing the boundaries of deep learning and computer vision capabilities.
B.3.2 Attack Perspective
From the attacker’s viewpoint, the effectiveness of physical attacks on deep learning-based vehicle detection systems is highly variant. Certain methodologies, such as ACTIVE, achieve astonishingly high success rates in defeating the detectors, with ASR values surpassing 70%. However, the majority of current attacks struggle to maintain comparable performance, often failing to reach even 20% ASR. This discrepancy can be partly attributed to the rapid advancements in detection algorithms, with the latest state-of-the-art models like EfficientNet, YOLO series, and RTMDet demonstrating increased resilience against known attacks. This disparity in the evolutionary pace between attackers and defenders underscores the importance of continuous research and innovation in adversarial attacks to keep pace with the evolving landscape of detection techniques.
Moreover, this evolving dynamic underscores a critical need for a more dynamic and collaborative ecosystem in deep learning and computer vision research. By closing the gap between attack methods and detector capabilities, the field will likely see increased robustness and security measures, ultimately benefiting automotive safety and other real-world applications relying on these systems.
On the other hand, when evaluating person detection, the outcome of physical attacks exhibits a different pattern, with ASR values typically remaining below 20%, and often even below 0%, which indicates that the attack method is less effective than random guessing and the eye-catching perturbation may arouse more attention than the object itself. Additionally, the variable transferability of these attack methodologies across different detectors leads to a wide disparity in ASR values. In certain instances, this manifests as negative ASR figures, indicative of a backfiring effect where the detectors become more adept at identifying targets in the presence of attempted attacks.
The significantly lower effectiveness of these attacks on pedestrian detection models highlights the comparative advantages of their 2D nature against primarily 2D adversarial perturbations. Nevertheless, the AdvTexture method, despite being a 2D approach, manages to incorporate 3D considerations, achieving higher ASRs compared to other attacks. This underscores the pivotal role of incorporating 3D awareness into attack strategies to exploit the vulnerabilities of pedestrian detectors more effectively.
These contrasting observations highlight the need for more sophisticated attack methods in the domain of pedestrian detection. By advancing the understanding of how 2D techniques can be adapted or combined with 3D concepts, attackers can create more potent adversarial samples, driving defender-side innovation to fortify models further. Such advancements will ultimately contribute to the progression of the field by promoting the design of more secure and reliable computer vision systems, particularly relevant in surveillance, autonomous navigation, and smart city infrastructures.
In summary, the diverse outcomes of physical attacks on both vehicle and person detection emphasize the importance of ongoing research and competition between attack and defense approaches. As the attacks become more intricate and align with the complex nature of real-world scenarios, deep learning and computer vision models will adapt, increasing their resilience and overall functionality. This continuous push-and-pull between adversaries and protectors fosters the evolution of robust, secure, and accurate object-detection technologies essential for numerous applications, including automotive safety, surveillance, and urban automation.
B.4 Additional overall experimental results
Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 4, 5, 6, 7, 8, 9, 10, and 11. In addition, the visualized evaluation results are shown in Fig. 9, 10, 11, 12, 13, 14, and 15.
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.83 | 0.477 | 0.231 | 0.545 | 0.606 | 0.502 | 0.434 | 0.678 | 0.235 | 0.532 |
AutoAssign | 0.786 | 0.574 | 0.37 | 0.559 | 0.589 | 0.609 | 0.415 | 0.722 | 0.487 | 0.648 |
CenterNet | 0.839 | 0.558 | 0.297 | 0.552 | 0.57 | 0.58 | 0.426 | 0.742 | 0.412 | 0.521 |
CentripetalNet | 0.78 | 0.685 | 0.558 | 0.725 | 0.687 | 0.725 | 0.527 | 0.801 | 0.501 | 0.648 |
CornerNet | 0.748 | 0.586 | 0.438 | 0.652 | 0.593 | 0.653 | 0.458 | 0.779 | 0.429 | 0.582 |
DDOD | 0.838 | 0.708 | 0.433 | 0.695 | 0.686 | 0.745 | 0.548 | 0.785 | 0.694 | 0.646 |
DyHead | 0.876 | 0.611 | 0.385 | 0.566 | 0.725 | 0.614 | 0.574 | 0.671 | 0.402 | 0.73 |
EfficientNet | 0.881 | 0.711 | 0.506 | 0.687 | 0.721 | 0.764 | 0.638 | 0.763 | 0.71 | 0.665 |
FCOS | 0.933 | 0.804 | 0.676 | 0.838 | 0.795 | 0.855 | 0.658 | 0.894 | 0.76 | 0.824 |
FoveaBox | 0.814 | 0.597 | 0.294 | 0.514 | 0.645 | 0.548 | 0.467 | 0.649 | 0.469 | 0.618 |
FreeAnchor | 0.81 | 0.51 | 0.381 | 0.611 | 0.563 | 0.643 | 0.431 | 0.638 | 0.336 | 0.582 |
FSAF | 0.788 | 0.51 | 0.233 | 0.566 | 0.559 | 0.537 | 0.432 | 0.661 | 0.364 | 0.529 |
GFL | 0.852 | 0.509 | 0.202 | 0.456 | 0.626 | 0.485 | 0.485 | 0.63 | 0.201 | 0.602 |
LD | 0.825 | 0.554 | 0.305 | 0.563 | 0.658 | 0.548 | 0.463 | 0.664 | 0.29 | 0.591 |
NAS-FPN | 0.87 | 0.623 | 0.473 | 0.662 | 0.673 | 0.695 | 0.5 | 0.764 | 0.382 | 0.655 |
PAA | 0.808 | 0.582 | 0.474 | 0.621 | 0.605 | 0.619 | 0.501 | 0.685 | 0.567 | 0.64 |
RetinaNet | 0.85 | 0.511 | 0.349 | 0.565 | 0.653 | 0.568 | 0.479 | 0.684 | 0.43 | 0.584 |
RTMDet | 0.875 | 0.717 | 0.625 | 0.771 | 0.733 | 0.753 | 0.736 | 0.821 | 0.676 | 0.72 |
TOOD | 0.781 | 0.495 | 0.353 | 0.522 | 0.584 | 0.557 | 0.462 | 0.615 | 0.37 | 0.572 |
VarifocalNet | 0.874 | 0.472 | 0.205 | 0.424 | 0.628 | 0.529 | 0.419 | 0.573 | 0.208 | 0.538 |
YOLOv5 | 0.886 | 0.76 | 0.744 | 0.762 | 0.807 | 0.821 | 0.788 | 0.857 | 0.812 | 0.75 |
YOLOv6 | 0.907 | 0.824 | 0.747 | 0.834 | 0.833 | 0.887 | 0.776 | 0.896 | 0.851 | 0.784 |
YOLOv7 | 0.906 | 0.774 | 0.762 | 0.829 | 0.822 | 0.866 | 0.786 | 0.903 | 0.774 | 0.803 |
YOLOv8 | 0.929 | 0.791 | 0.761 | 0.812 | 0.838 | 0.873 | 0.793 | 0.917 | 0.74 | 0.803 |
YOLOX | 0.908 | 0.766 | 0.683 | 0.783 | 0.817 | 0.851 | 0.794 | 0.849 | 0.702 | 0.782 |
Faster R-CNN | 0.772 | 0.375 | 0.141 | 0.338 | 0.509 | 0.46 | 0.369 | 0.612 | 0.268 | 0.438 |
Cascade R-CNN | 0.802 | 0.483 | 0.297 | 0.488 | 0.574 | 0.607 | 0.407 | 0.673 | 0.334 | 0.532 |
Cascade RPN | 0.805 | 0.53 | 0.291 | 0.452 | 0.588 | 0.55 | 0.441 | 0.662 | 0.452 | 0.548 |
Double Heads | 0.797 | 0.521 | 0.295 | 0.539 | 0.537 | 0.621 | 0.404 | 0.713 | 0.445 | 0.516 |
FPG | 0.846 | 0.678 | 0.486 | 0.714 | 0.671 | 0.749 | 0.503 | 0.806 | 0.47 | 0.65 |
Grid R-CNN | 0.795 | 0.472 | 0.244 | 0.513 | 0.529 | 0.65 | 0.399 | 0.699 | 0.392 | 0.494 |
Guided Anchoring | 0.904 | 0.723 | 0.555 | 0.747 | 0.748 | 0.781 | 0.524 | 0.828 | 0.738 | 0.717 |
HRNet | 0.76 | 0.547 | 0.29 | 0.52 | 0.512 | 0.571 | 0.336 | 0.738 | 0.462 | 0.511 |
Libra R-CNN | 0.78 | 0.49 | 0.294 | 0.527 | 0.563 | 0.492 | 0.486 | 0.664 | 0.334 | 0.54 |
PAFPN | 0.8 | 0.497 | 0.206 | 0.463 | 0.562 | 0.54 | 0.413 | 0.682 | 0.383 | 0.457 |
RepPoints | 0.847 | 0.576 | 0.222 | 0.525 | 0.547 | 0.557 | 0.427 | 0.712 | 0.565 | 0.523 |
Res2Net | 0.874 | 0.64 | 0.494 | 0.698 | 0.72 | 0.74 | 0.482 | 0.797 | 0.601 | 0.716 |
ResNeSt | 0.837 | 0.502 | 0.352 | 0.535 | 0.587 | 0.499 | 0.538 | 0.493 | 0.407 | 0.555 |
SABL | 0.796 | 0.46 | 0.262 | 0.501 | 0.563 | 0.535 | 0.423 | 0.648 | 0.359 | 0.484 |
Sparse R-CNN | 0.774 | 0.469 | 0.257 | 0.398 | 0.604 | 0.418 | 0.44 | 0.518 | 0.316 | 0.532 |
DETR | 0.636 | 0.114 | 0.048 | 0.033 | 0.351 | 0.17 | 0.333 | 0.198 | 0.025 | 0.339 |
Conditional DETR | 0.793 | 0.554 | 0.408 | 0.575 | 0.644 | 0.617 | 0.525 | 0.7 | 0.457 | 0.671 |
DDQ | 0.809 | 0.55 | 0.457 | 0.531 | 0.649 | 0.676 | 0.631 | 0.626 | 0.453 | 0.629 |
DAB-DETR | 0.838 | 0.391 | 0.194 | 0.308 | 0.616 | 0.473 | 0.488 | 0.576 | 0.163 | 0.526 |
Deformable DETR | 0.827 | 0.642 | 0.371 | 0.528 | 0.641 | 0.67 | 0.46 | 0.662 | 0.525 | 0.626 |
DINO | 0.78 | 0.351 | 0.236 | 0.322 | 0.543 | 0.56 | 0.423 | 0.526 | 0.217 | 0.49 |
PVT | 0.828 | 0.719 | 0.355 | 0.648 | 0.711 | 0.802 | 0.547 | 0.853 | 0.592 | 0.52 |
PVTv2 | 0.845 | 0.666 | 0.494 | 0.763 | 0.621 | 0.844 | 0.425 | 0.803 | 0.704 | 0.476 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.238 | 0.156 | 0.077 | 0.182 | 0.183 | 0.164 | 0.133 | 0.212 | 0.087 | 0.166 |
AutoAssign | 0.238 | 0.183 | 0.126 | 0.189 | 0.182 | 0.199 | 0.139 | 0.236 | 0.164 | 0.212 |
CenterNet | 0.238 | 0.167 | 0.093 | 0.178 | 0.165 | 0.182 | 0.14 | 0.23 | 0.126 | 0.156 |
CentripetalNet | 0.228 | 0.215 | 0.164 | 0.225 | 0.206 | 0.22 | 0.161 | 0.245 | 0.155 | 0.194 |
CornerNet | 0.218 | 0.184 | 0.132 | 0.198 | 0.175 | 0.199 | 0.142 | 0.237 | 0.129 | 0.173 |
DDOD | 0.242 | 0.21 | 0.127 | 0.221 | 0.202 | 0.234 | 0.174 | 0.242 | 0.211 | 0.193 |
DyHead | 0.26 | 0.196 | 0.121 | 0.18 | 0.221 | 0.191 | 0.18 | 0.21 | 0.126 | 0.227 |
EfficientNet | 0.252 | 0.225 | 0.168 | 0.217 | 0.219 | 0.239 | 0.206 | 0.243 | 0.232 | 0.21 |
FCOS | 0.277 | 0.251 | 0.211 | 0.266 | 0.246 | 0.272 | 0.214 | 0.285 | 0.236 | 0.256 |
FoveaBox | 0.235 | 0.184 | 0.089 | 0.165 | 0.191 | 0.17 | 0.145 | 0.196 | 0.149 | 0.193 |
FreeAnchor | 0.241 | 0.159 | 0.111 | 0.19 | 0.179 | 0.205 | 0.138 | 0.205 | 0.098 | 0.184 |
FSAF | 0.231 | 0.171 | 0.077 | 0.187 | 0.181 | 0.181 | 0.141 | 0.213 | 0.129 | 0.179 |
GFL | 0.244 | 0.169 | 0.064 | 0.152 | 0.192 | 0.163 | 0.157 | 0.201 | 0.069 | 0.192 |
LD | 0.235 | 0.176 | 0.095 | 0.181 | 0.205 | 0.176 | 0.146 | 0.208 | 0.094 | 0.187 |
NAS-FPN | 0.256 | 0.199 | 0.146 | 0.209 | 0.213 | 0.215 | 0.16 | 0.251 | 0.124 | 0.208 |
PAA | 0.237 | 0.178 | 0.139 | 0.189 | 0.178 | 0.195 | 0.157 | 0.208 | 0.173 | 0.194 |
RetinaNet | 0.249 | 0.169 | 0.108 | 0.194 | 0.209 | 0.189 | 0.166 | 0.227 | 0.155 | 0.191 |
RTMDet | 0.254 | 0.227 | 0.185 | 0.235 | 0.224 | 0.232 | 0.239 | 0.241 | 0.209 | 0.221 |
TOOD | 0.231 | 0.155 | 0.109 | 0.159 | 0.173 | 0.176 | 0.14 | 0.183 | 0.118 | 0.169 |
VarifocalNet | 0.248 | 0.144 | 0.063 | 0.127 | 0.188 | 0.164 | 0.132 | 0.175 | 0.062 | 0.162 |
YOLOv5 | 0.259 | 0.227 | 0.223 | 0.23 | 0.237 | 0.249 | 0.244 | 0.253 | 0.246 | 0.221 |
YOLOv6 | 0.256 | 0.245 | 0.218 | 0.251 | 0.242 | 0.262 | 0.238 | 0.263 | 0.25 | 0.229 |
YOLOv7 | 0.265 | 0.241 | 0.225 | 0.252 | 0.246 | 0.262 | 0.241 | 0.272 | 0.227 | 0.243 |
YOLOv8 | 0.276 | 0.246 | 0.239 | 0.254 | 0.252 | 0.269 | 0.256 | 0.283 | 0.236 | 0.246 |
YOLOX | 0.263 | 0.233 | 0.212 | 0.236 | 0.237 | 0.253 | 0.248 | 0.251 | 0.217 | 0.233 |
Faster R-CNN | 0.212 | 0.117 | 0.042 | 0.11 | 0.159 | 0.145 | 0.111 | 0.193 | 0.087 | 0.137 |
Cascade R-CNN | 0.232 | 0.152 | 0.088 | 0.15 | 0.182 | 0.19 | 0.135 | 0.214 | 0.115 | 0.169 |
Cascade RPN | 0.229 | 0.157 | 0.083 | 0.132 | 0.175 | 0.157 | 0.137 | 0.201 | 0.122 | 0.166 |
Double Heads | 0.238 | 0.168 | 0.09 | 0.179 | 0.171 | 0.205 | 0.143 | 0.231 | 0.15 | 0.164 |
FPG | 0.247 | 0.222 | 0.149 | 0.224 | 0.21 | 0.233 | 0.161 | 0.265 | 0.152 | 0.216 |
Grid R-CNN | 0.231 | 0.154 | 0.078 | 0.173 | 0.17 | 0.208 | 0.135 | 0.225 | 0.127 | 0.164 |
Guided Anchoring | 0.269 | 0.239 | 0.176 | 0.244 | 0.235 | 0.243 | 0.174 | 0.265 | 0.244 | 0.229 |
HRNet | 0.219 | 0.171 | 0.091 | 0.164 | 0.159 | 0.173 | 0.114 | 0.236 | 0.148 | 0.165 |
Libra R-CNN | 0.234 | 0.161 | 0.094 | 0.18 | 0.179 | 0.169 | 0.159 | 0.222 | 0.115 | 0.175 |
PAFPN | 0.219 | 0.147 | 0.057 | 0.145 | 0.17 | 0.168 | 0.122 | 0.205 | 0.129 | 0.139 |
RepPoints | 0.251 | 0.18 | 0.068 | 0.167 | 0.172 | 0.185 | 0.141 | 0.234 | 0.189 | 0.166 |
Res2Net | 0.25 | 0.195 | 0.154 | 0.217 | 0.212 | 0.22 | 0.162 | 0.244 | 0.192 | 0.22 |
ResNeSt | 0.23 | 0.156 | 0.101 | 0.158 | 0.174 | 0.154 | 0.169 | 0.142 | 0.126 | 0.162 |
SABL | 0.233 | 0.146 | 0.08 | 0.159 | 0.173 | 0.177 | 0.139 | 0.199 | 0.123 | 0.155 |
Sparse R-CNN | 0.238 | 0.159 | 0.095 | 0.137 | 0.195 | 0.145 | 0.154 | 0.172 | 0.128 | 0.168 |
DETR | 0.186 | 0.047 | 0.017 | 0.01 | 0.113 | 0.059 | 0.111 | 0.065 | 0.009 | 0.105 |
Conditional DETR | 0.236 | 0.183 | 0.129 | 0.183 | 0.199 | 0.205 | 0.172 | 0.211 | 0.154 | 0.212 |
DDQ | 0.232 | 0.164 | 0.133 | 0.152 | 0.185 | 0.2 | 0.189 | 0.182 | 0.136 | 0.187 |
DAB-DETR | 0.239 | 0.115 | 0.057 | 0.09 | 0.178 | 0.145 | 0.143 | 0.157 | 0.05 | 0.143 |
Deformable DETR | 0.229 | 0.187 | 0.106 | 0.147 | 0.177 | 0.197 | 0.136 | 0.188 | 0.158 | 0.171 |
DINO | 0.232 | 0.11 | 0.075 | 0.104 | 0.172 | 0.18 | 0.139 | 0.161 | 0.078 | 0.156 |
PVT | 0.229 | 0.229 | 0.106 | 0.213 | 0.224 | 0.254 | 0.177 | 0.26 | 0.199 | 0.163 |
PVTv2 | 0.24 | 0.214 | 0.154 | 0.252 | 0.195 | 0.275 | 0.147 | 0.244 | 0.237 | 0.149 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.973 | 0.808 | 0.518 | 0.84 | 0.883 | 0.811 | 0.732 | 0.924 | 0.549 | 0.826 |
AutoAssign | 0.946 | 0.846 | 0.703 | 0.876 | 0.849 | 0.888 | 0.763 | 0.938 | 0.856 | 0.901 |
CenterNet | 0.976 | 0.926 | 0.674 | 0.893 | 0.901 | 0.897 | 0.806 | 0.97 | 0.873 | 0.895 |
CentripetalNet | 0.943 | 0.867 | 0.753 | 0.908 | 0.901 | 0.918 | 0.749 | 0.958 | 0.81 | 0.853 |
CornerNet | 0.948 | 0.8 | 0.692 | 0.905 | 0.847 | 0.902 | 0.697 | 0.959 | 0.722 | 0.829 |
DDOD | 0.95 | 0.936 | 0.756 | 0.923 | 0.9 | 0.934 | 0.863 | 0.948 | 0.949 | 0.909 |
DyHead | 0.977 | 0.836 | 0.606 | 0.792 | 0.903 | 0.845 | 0.838 | 0.843 | 0.685 | 0.927 |
EfficientNet | 0.977 | 0.974 | 0.912 | 0.975 | 0.974 | 0.97 | 0.951 | 0.973 | 0.966 | 0.948 |
FCOS | 0.981 | 0.974 | 0.93 | 0.968 | 0.951 | 0.975 | 0.931 | 0.972 | 0.939 | 0.959 |
FoveaBox | 0.955 | 0.873 | 0.623 | 0.806 | 0.896 | 0.832 | 0.748 | 0.881 | 0.768 | 0.878 |
FreeAnchor | 0.973 | 0.841 | 0.855 | 0.928 | 0.864 | 0.939 | 0.781 | 0.89 | 0.68 | 0.887 |
FSAF | 0.95 | 0.836 | 0.567 | 0.886 | 0.823 | 0.84 | 0.632 | 0.918 | 0.659 | 0.8 |
GFL | 0.978 | 0.837 | 0.575 | 0.81 | 0.908 | 0.839 | 0.836 | 0.913 | 0.549 | 0.903 |
LD | 0.98 | 0.893 | 0.579 | 0.862 | 0.927 | 0.871 | 0.743 | 0.917 | 0.676 | 0.91 |
NAS-FPN | 0.975 | 0.916 | 0.768 | 0.932 | 0.944 | 0.946 | 0.817 | 0.937 | 0.766 | 0.925 |
PAA | 0.966 | 0.923 | 0.861 | 0.952 | 0.905 | 0.937 | 0.895 | 0.938 | 0.951 | 0.92 |
RetinaNet | 0.98 | 0.859 | 0.764 | 0.915 | 0.934 | 0.89 | 0.775 | 0.921 | 0.729 | 0.9 |
RTMDet | 0.982 | 0.954 | 0.943 | 0.971 | 0.958 | 0.971 | 0.975 | 0.98 | 0.99 | 0.937 |
TOOD | 0.908 | 0.763 | 0.666 | 0.83 | 0.834 | 0.826 | 0.717 | 0.847 | 0.622 | 0.853 |
VarifocalNet | 0.977 | 0.802 | 0.486 | 0.77 | 0.9 | 0.832 | 0.671 | 0.866 | 0.466 | 0.829 |
YOLOv5 | 0.975 | 0.979 | 0.967 | 0.974 | 0.977 | 0.974 | 0.974 | 0.972 | 0.985 | 0.966 |
YOLOv6 | 0.985 | 0.982 | 0.968 | 0.974 | 0.983 | 0.979 | 0.974 | 0.984 | 0.978 | 0.973 |
YOLOv7 | 0.962 | 0.924 | 0.931 | 0.948 | 0.939 | 0.958 | 0.932 | 0.96 | 0.932 | 0.931 |
YOLOv8 | 0.975 | 0.919 | 0.903 | 0.93 | 0.942 | 0.96 | 0.916 | 0.965 | 0.885 | 0.917 |
YOLOX | 0.955 | 0.877 | 0.853 | 0.902 | 0.91 | 0.945 | 0.926 | 0.918 | 0.868 | 0.891 |
Faster R-CNN | 0.846 | 0.479 | 0.268 | 0.493 | 0.595 | 0.593 | 0.417 | 0.752 | 0.337 | 0.532 |
Cascade R-CNN | 0.854 | 0.539 | 0.382 | 0.591 | 0.65 | 0.689 | 0.448 | 0.78 | 0.368 | 0.602 |
Cascade RPN | 0.973 | 0.884 | 0.741 | 0.933 | 0.898 | 0.957 | 0.885 | 0.967 | 0.868 | 0.891 |
Double Heads | 0.849 | 0.597 | 0.416 | 0.654 | 0.621 | 0.725 | 0.459 | 0.819 | 0.488 | 0.594 |
FPG | 0.912 | 0.763 | 0.624 | 0.848 | 0.761 | 0.866 | 0.587 | 0.907 | 0.61 | 0.748 |
Grid R-CNN | 0.873 | 0.585 | 0.39 | 0.672 | 0.653 | 0.794 | 0.488 | 0.836 | 0.495 | 0.614 |
Guided Anchoring | 0.975 | 0.962 | 0.914 | 0.966 | 0.962 | 0.966 | 0.839 | 0.961 | 0.929 | 0.956 |
HRNet | 0.844 | 0.655 | 0.429 | 0.687 | 0.627 | 0.732 | 0.423 | 0.875 | 0.532 | 0.609 |
Libra R-CNN | 0.959 | 0.828 | 0.599 | 0.824 | 0.892 | 0.775 | 0.805 | 0.92 | 0.563 | 0.858 |
PAFPN | 0.856 | 0.61 | 0.337 | 0.591 | 0.661 | 0.659 | 0.49 | 0.805 | 0.434 | 0.569 |
RepPoints | 0.978 | 0.903 | 0.595 | 0.874 | 0.885 | 0.884 | 0.833 | 0.945 | 0.883 | 0.848 |
Res2Net | 0.911 | 0.692 | 0.541 | 0.757 | 0.763 | 0.793 | 0.511 | 0.852 | 0.646 | 0.761 |
ResNeSt | 0.929 | 0.646 | 0.497 | 0.727 | 0.701 | 0.691 | 0.685 | 0.802 | 0.515 | 0.716 |
SABL | 0.856 | 0.545 | 0.387 | 0.646 | 0.656 | 0.636 | 0.488 | 0.774 | 0.407 | 0.571 |
Sparse R-CNN | 0.959 | 0.877 | 0.513 | 0.759 | 0.913 | 0.746 | 0.733 | 0.882 | 0.605 | 0.871 |
DETR | 0.746 | 0.328 | 0.232 | 0.256 | 0.468 | 0.383 | 0.457 | 0.369 | 0.234 | 0.468 |
Conditional DETR | 0.962 | 0.831 | 0.73 | 0.931 | 0.865 | 0.934 | 0.881 | 0.964 | 0.839 | 0.917 |
DDQ | 0.983 | 0.976 | 0.972 | 0.979 | 0.975 | 0.975 | 0.977 | 0.974 | 0.983 | 0.969 |
DAB-DETR | 0.98 | 0.909 | 0.928 | 0.97 | 0.948 | 0.968 | 0.946 | 0.924 | 0.91 | 0.934 |
Deformable DETR | 0.954 | 0.907 | 0.704 | 0.902 | 0.879 | 0.924 | 0.766 | 0.905 | 0.91 | 0.88 |
DINO | 0.975 | 0.895 | 0.883 | 0.953 | 0.923 | 0.958 | 0.922 | 0.953 | 0.912 | 0.924 |
PVT | 0.948 | 0.953 | 0.827 | 0.936 | 0.957 | 0.956 | 0.901 | 0.963 | 0.954 | 0.886 |
PVTv2 | 0.973 | 0.942 | 0.884 | 0.952 | 0.934 | 0.973 | 0.835 | 0.967 | 0.941 | 0.867 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.374 | 0.318 | 0.206 | 0.341 | 0.342 | 0.321 | 0.296 | 0.371 | 0.219 | 0.325 |
AutoAssign | 0.385 | 0.341 | 0.293 | 0.366 | 0.335 | 0.367 | 0.32 | 0.39 | 0.345 | 0.374 |
CenterNet | 0.396 | 0.365 | 0.275 | 0.373 | 0.353 | 0.373 | 0.342 | 0.408 | 0.344 | 0.361 |
CentripetalNet | 0.378 | 0.362 | 0.313 | 0.386 | 0.365 | 0.384 | 0.31 | 0.404 | 0.321 | 0.35 |
CornerNet | 0.387 | 0.337 | 0.286 | 0.382 | 0.343 | 0.38 | 0.292 | 0.401 | 0.282 | 0.336 |
DDOD | 0.366 | 0.363 | 0.302 | 0.371 | 0.347 | 0.379 | 0.357 | 0.383 | 0.366 | 0.358 |
DyHead | 0.378 | 0.338 | 0.254 | 0.32 | 0.35 | 0.338 | 0.342 | 0.336 | 0.255 | 0.364 |
EfficientNet | 0.387 | 0.397 | 0.377 | 0.395 | 0.393 | 0.394 | 0.399 | 0.402 | 0.406 | 0.38 |
FCOS | 0.401 | 0.4 | 0.382 | 0.399 | 0.389 | 0.405 | 0.389 | 0.418 | 0.37 | 0.397 |
FoveaBox | 0.371 | 0.337 | 0.246 | 0.327 | 0.344 | 0.333 | 0.304 | 0.353 | 0.303 | 0.344 |
FreeAnchor | 0.384 | 0.332 | 0.333 | 0.367 | 0.345 | 0.38 | 0.324 | 0.365 | 0.244 | 0.357 |
FSAF | 0.37 | 0.337 | 0.229 | 0.362 | 0.33 | 0.342 | 0.271 | 0.377 | 0.257 | 0.325 |
GFL | 0.375 | 0.337 | 0.238 | 0.336 | 0.354 | 0.341 | 0.348 | 0.373 | 0.226 | 0.357 |
LD | 0.372 | 0.352 | 0.236 | 0.351 | 0.367 | 0.348 | 0.314 | 0.367 | 0.268 | 0.36 |
NAS-FPN | 0.38 | 0.367 | 0.307 | 0.373 | 0.378 | 0.375 | 0.34 | 0.384 | 0.296 | 0.373 |
PAA | 0.399 | 0.371 | 0.337 | 0.385 | 0.357 | 0.39 | 0.376 | 0.388 | 0.383 | 0.369 |
RetinaNet | 0.392 | 0.343 | 0.298 | 0.371 | 0.378 | 0.367 | 0.34 | 0.39 | 0.301 | 0.364 |
RTMDet | 0.357 | 0.322 | 0.293 | 0.345 | 0.319 | 0.351 | 0.34 | 0.362 | 0.304 | 0.311 |
TOOD | 0.345 | 0.293 | 0.261 | 0.332 | 0.312 | 0.329 | 0.283 | 0.327 | 0.236 | 0.323 |
VarifocalNet | 0.376 | 0.31 | 0.188 | 0.304 | 0.352 | 0.329 | 0.277 | 0.338 | 0.173 | 0.321 |
YOLOv5 | 0.364 | 0.364 | 0.366 | 0.371 | 0.358 | 0.378 | 0.377 | 0.373 | 0.354 | 0.355 |
YOLOv6 | 0.357 | 0.361 | 0.343 | 0.363 | 0.352 | 0.37 | 0.359 | 0.37 | 0.351 | 0.343 |
YOLOv7 | 0.366 | 0.356 | 0.339 | 0.36 | 0.355 | 0.371 | 0.358 | 0.376 | 0.338 | 0.352 |
YOLOv8 | 0.376 | 0.358 | 0.354 | 0.369 | 0.357 | 0.376 | 0.368 | 0.385 | 0.336 | 0.354 |
YOLOX | 0.362 | 0.33 | 0.325 | 0.345 | 0.338 | 0.368 | 0.367 | 0.358 | 0.309 | 0.332 |
Faster R-CNN | 0.303 | 0.187 | 0.099 | 0.194 | 0.228 | 0.236 | 0.163 | 0.302 | 0.134 | 0.203 |
Cascade R-CNN | 0.316 | 0.213 | 0.147 | 0.232 | 0.255 | 0.272 | 0.185 | 0.312 | 0.145 | 0.239 |
Cascade RPN | 0.375 | 0.333 | 0.276 | 0.351 | 0.341 | 0.358 | 0.353 | 0.376 | 0.307 | 0.341 |
Double Heads | 0.32 | 0.239 | 0.165 | 0.27 | 0.249 | 0.299 | 0.2 | 0.333 | 0.193 | 0.238 |
FPG | 0.347 | 0.31 | 0.249 | 0.35 | 0.302 | 0.352 | 0.239 | 0.379 | 0.249 | 0.306 |
Grid R-CNN | 0.326 | 0.23 | 0.155 | 0.273 | 0.259 | 0.312 | 0.205 | 0.33 | 0.188 | 0.244 |
Guided Anchoring | 0.38 | 0.396 | 0.373 | 0.39 | 0.385 | 0.384 | 0.35 | 0.399 | 0.363 | 0.379 |
HRNet | 0.303 | 0.239 | 0.159 | 0.252 | 0.228 | 0.264 | 0.167 | 0.333 | 0.188 | 0.226 |
Libra R-CNN | 0.383 | 0.326 | 0.233 | 0.334 | 0.357 | 0.315 | 0.333 | 0.385 | 0.223 | 0.337 |
PAFPN | 0.304 | 0.226 | 0.127 | 0.228 | 0.249 | 0.254 | 0.185 | 0.304 | 0.167 | 0.214 |
RepPoints | 0.387 | 0.356 | 0.235 | 0.355 | 0.349 | 0.372 | 0.358 | 0.393 | 0.364 | 0.336 |
Res2Net | 0.341 | 0.261 | 0.208 | 0.302 | 0.284 | 0.305 | 0.21 | 0.338 | 0.237 | 0.292 |
ResNeSt | 0.341 | 0.257 | 0.189 | 0.274 | 0.261 | 0.27 | 0.268 | 0.302 | 0.201 | 0.271 |
SABL | 0.318 | 0.213 | 0.151 | 0.26 | 0.253 | 0.257 | 0.204 | 0.303 | 0.16 | 0.226 |
Sparse R-CNN | 0.395 | 0.367 | 0.219 | 0.324 | 0.38 | 0.32 | 0.312 | 0.381 | 0.249 | 0.361 |
DETR | 0.334 | 0.144 | 0.105 | 0.106 | 0.203 | 0.171 | 0.204 | 0.158 | 0.085 | 0.203 |
Conditional DETR | 0.387 | 0.341 | 0.32 | 0.404 | 0.342 | 0.406 | 0.367 | 0.391 | 0.328 | 0.374 |
DDQ | 0.393 | 0.389 | 0.388 | 0.394 | 0.39 | 0.389 | 0.403 | 0.391 | 0.374 | 0.389 |
DAB-DETR | 0.404 | 0.366 | 0.384 | 0.406 | 0.385 | 0.413 | 0.404 | 0.376 | 0.36 | 0.383 |
Deformable DETR | 0.378 | 0.352 | 0.277 | 0.351 | 0.333 | 0.372 | 0.3 | 0.351 | 0.339 | 0.34 |
DINO | 0.388 | 0.353 | 0.353 | 0.39 | 0.364 | 0.388 | 0.372 | 0.384 | 0.346 | 0.37 |
PVT | 0.349 | 0.371 | 0.319 | 0.365 | 0.37 | 0.379 | 0.365 | 0.382 | 0.387 | 0.34 |
PVTv2 | 0.372 | 0.372 | 0.359 | 0.386 | 0.365 | 0.399 | 0.354 | 0.39 | 0.384 | 0.343 |
Clean |
Random |
AdvCam |
UPC |
NatPatch |
MTD |
LAP |
InvisCloak |
DAP |
AdvTshirt |
AdvTexture |
AdvPatch |
AdvPattern |
AdvCaT |
ATSS | 0.54 | 0.517 | 0.498 | 0.428 | 0.419 | 0.473 | 0.522 | 0.468 | 0.495 | 0.458 | 0.385 | 0.454 | 0.492 | 0.514 |
AutoAssign | 0.491 | 0.466 | 0.454 | 0.314 | 0.36 | 0.423 | 0.456 | 0.43 | 0.403 | 0.41 | 0.346 | 0.427 | 0.453 | 0.484 |
CenterNet | 0.524 | 0.476 | 0.477 | 0.408 | 0.39 | 0.469 | 0.483 | 0.436 | 0.437 | 0.45 | 0.372 | 0.43 | 0.474 | 0.524 |
CentripetalNet | 0.526 | 0.53 | 0.524 | 0.48 | 0.349 | 0.524 | 0.508 | 0.51 | 0.471 | 0.473 | 0.405 | 0.472 | 0.515 | 0.526 |
CornerNet | 0.517 | 0.51 | 0.505 | 0.403 | 0.295 | 0.488 | 0.494 | 0.449 | 0.444 | 0.42 | 0.345 | 0.414 | 0.48 | 0.506 |
DDOD | 0.481 | 0.48 | 0.448 | 0.359 | 0.416 | 0.421 | 0.47 | 0.445 | 0.453 | 0.433 | 0.329 | 0.409 | 0.442 | 0.45 |
DyHead | 0.474 | 0.483 | 0.485 | 0.406 | 0.402 | 0.464 | 0.501 | 0.454 | 0.473 | 0.433 | 0.4 | 0.433 | 0.467 | 0.474 |
EfficientNet | 0.457 | 0.431 | 0.442 | 0.398 | 0.418 | 0.406 | 0.431 | 0.399 | 0.407 | 0.403 | 0.394 | 0.396 | 0.422 | 0.433 |
FCOS | 0.45 | 0.438 | 0.429 | 0.364 | 0.383 | 0.41 | 0.433 | 0.407 | 0.404 | 0.407 | 0.356 | 0.409 | 0.425 | 0.448 |
FoveaBox | 0.543 | 0.53 | 0.536 | 0.473 | 0.475 | 0.482 | 0.54 | 0.481 | 0.523 | 0.483 | 0.374 | 0.458 | 0.498 | 0.533 |
FreeAnchor | 0.537 | 0.522 | 0.493 | 0.414 | 0.396 | 0.431 | 0.492 | 0.443 | 0.446 | 0.411 | 0.331 | 0.447 | 0.48 | 0.513 |
FSAF | 0.554 | 0.551 | 0.529 | 0.444 | 0.439 | 0.485 | 0.538 | 0.496 | 0.515 | 0.457 | 0.379 | 0.479 | 0.512 | 0.527 |
GFL | 0.57 | 0.541 | 0.532 | 0.431 | 0.453 | 0.495 | 0.509 | 0.478 | 0.495 | 0.48 | 0.398 | 0.459 | 0.52 | 0.547 |
LD | 0.57 | 0.54 | 0.524 | 0.401 | 0.397 | 0.486 | 0.517 | 0.484 | 0.489 | 0.477 | 0.385 | 0.484 | 0.519 | 0.535 |
NAS-FPN | 0.442 | 0.436 | 0.433 | 0.365 | 0.391 | 0.4 | 0.451 | 0.399 | 0.394 | 0.398 | 0.32 | 0.399 | 0.413 | 0.435 |
PAA | 0.464 | 0.464 | 0.457 | 0.402 | 0.389 | 0.43 | 0.451 | 0.432 | 0.447 | 0.402 | 0.322 | 0.409 | 0.441 | 0.463 |
RetinaNet | 0.522 | 0.543 | 0.497 | 0.438 | 0.425 | 0.459 | 0.505 | 0.483 | 0.478 | 0.454 | 0.384 | 0.475 | 0.489 | 0.528 |
RTMDet | 0.533 | 0.482 | 0.515 | 0.52 | 0.466 | 0.46 | 0.495 | 0.437 | 0.472 | 0.47 | 0.459 | 0.449 | 0.466 | 0.5 |
TOOD | 0.474 | 0.5 | 0.475 | 0.384 | 0.376 | 0.453 | 0.503 | 0.453 | 0.486 | 0.441 | 0.371 | 0.435 | 0.457 | 0.475 |
VarifocalNet | 0.492 | 0.505 | 0.481 | 0.387 | 0.395 | 0.443 | 0.504 | 0.444 | 0.469 | 0.445 | 0.368 | 0.436 | 0.47 | 0.506 |
YOLOv5 | 0.481 | 0.46 | 0.472 | 0.448 | 0.403 | 0.453 | 0.484 | 0.454 | 0.485 | 0.418 | 0.35 | 0.42 | 0.452 | 0.459 |
YOLOv6 | 0.467 | 0.445 | 0.461 | 0.456 | 0.435 | 0.438 | 0.444 | 0.446 | 0.459 | 0.437 | 0.423 | 0.432 | 0.444 | 0.449 |
YOLOv7 | 0.463 | 0.438 | 0.48 | 0.446 | 0.343 | 0.401 | 0.438 | 0.403 | 0.457 | 0.402 | 0.372 | 0.366 | 0.429 | 0.437 |
YOLOv8 | 0.434 | 0.421 | 0.431 | 0.432 | 0.402 | 0.415 | 0.421 | 0.416 | 0.429 | 0.416 | 0.405 | 0.409 | 0.415 | 0.417 |
YOLOX | 0.448 | 0.436 | 0.457 | 0.457 | 0.382 | 0.432 | 0.46 | 0.425 | 0.46 | 0.433 | 0.393 | 0.412 | 0.426 | 0.437 |
Faster R-CNN | 0.541 | 0.547 | 0.497 | 0.416 | 0.425 | 0.456 | 0.532 | 0.468 | 0.456 | 0.448 | 0.341 | 0.432 | 0.512 | 0.534 |
Cascade R-CNN | 0.559 | 0.551 | 0.539 | 0.431 | 0.445 | 0.488 | 0.551 | 0.463 | 0.508 | 0.479 | 0.355 | 0.454 | 0.523 | 0.55 |
Cascade RPN | 0.538 | 0.537 | 0.528 | 0.389 | 0.407 | 0.482 | 0.508 | 0.472 | 0.483 | 0.461 | 0.335 | 0.445 | 0.508 | 0.532 |
Double Heads | 0.552 | 0.526 | 0.531 | 0.419 | 0.408 | 0.46 | 0.533 | 0.442 | 0.489 | 0.454 | 0.359 | 0.44 | 0.496 | 0.527 |
FPG | 0.462 | 0.473 | 0.451 | 0.413 | 0.395 | 0.415 | 0.466 | 0.408 | 0.424 | 0.405 | 0.317 | 0.409 | 0.444 | 0.45 |
Grid R-CNN | 0.512 | 0.502 | 0.492 | 0.397 | 0.404 | 0.449 | 0.517 | 0.462 | 0.471 | 0.43 | 0.363 | 0.424 | 0.481 | 0.488 |
Guided Anchoring | 0.497 | 0.537 | 0.504 | 0.427 | 0.396 | 0.47 | 0.525 | 0.479 | 0.454 | 0.449 | 0.375 | 0.452 | 0.494 | 0.502 |
HRNet | 0.498 | 0.489 | 0.495 | 0.457 | 0.404 | 0.453 | 0.489 | 0.47 | 0.442 | 0.472 | 0.419 | 0.437 | 0.443 | 0.497 |
Libra R-CNN | 0.535 | 0.517 | 0.479 | 0.452 | 0.404 | 0.446 | 0.468 | 0.433 | 0.447 | 0.431 | 0.374 | 0.421 | 0.442 | 0.494 |
PAFPN | 0.539 | 0.534 | 0.529 | 0.429 | 0.438 | 0.468 | 0.522 | 0.477 | 0.47 | 0.458 | 0.349 | 0.447 | 0.516 | 0.559 |
RepPoints | 0.572 | 0.559 | 0.53 | 0.434 | 0.475 | 0.478 | 0.535 | 0.47 | 0.504 | 0.455 | 0.38 | 0.451 | 0.525 | 0.56 |
Res2Net | 0.449 | 0.437 | 0.435 | 0.403 | 0.301 | 0.402 | 0.458 | 0.406 | 0.407 | 0.386 | 0.324 | 0.387 | 0.428 | 0.451 |
ResNeSt | 0.443 | 0.455 | 0.409 | 0.396 | 0.396 | 0.405 | 0.432 | 0.385 | 0.364 | 0.374 | 0.358 | 0.374 | 0.416 | 0.451 |
SABL | 0.563 | 0.559 | 0.525 | 0.418 | 0.471 | 0.491 | 0.534 | 0.503 | 0.498 | 0.496 | 0.382 | 0.484 | 0.534 | 0.552 |
Sparse R-CNN | 0.492 | 0.481 | 0.477 | 0.38 | 0.347 | 0.434 | 0.484 | 0.386 | 0.396 | 0.407 | 0.352 | 0.389 | 0.455 | 0.493 |
DETR | 0.553 | 0.497 | 0.481 | 0.337 | 0.318 | 0.467 | 0.49 | 0.466 | 0.432 | 0.473 | 0.343 | 0.444 | 0.475 | 0.51 |
Conditional DETR | 0.535 | 0.497 | 0.453 | 0.378 | 0.351 | 0.424 | 0.449 | 0.401 | 0.44 | 0.416 | 0.294 | 0.412 | 0.422 | 0.485 |
DDQ | 0.449 | 0.454 | 0.452 | 0.377 | 0.388 | 0.424 | 0.451 | 0.426 | 0.408 | 0.422 | 0.34 | 0.421 | 0.439 | 0.451 |
DAB-DETR | 0.441 | 0.429 | 0.428 | 0.344 | 0.36 | 0.371 | 0.407 | 0.357 | 0.373 | 0.374 | 0.276 | 0.336 | 0.392 | 0.44 |
Deformable DETR | 0.475 | 0.462 | 0.475 | 0.345 | 0.389 | 0.421 | 0.46 | 0.442 | 0.418 | 0.424 | 0.286 | 0.416 | 0.471 | 0.45 |
DINO | 0.419 | 0.421 | 0.416 | 0.337 | 0.283 | 0.385 | 0.415 | 0.402 | 0.375 | 0.391 | 0.316 | 0.378 | 0.394 | 0.423 |
PVT | 0.474 | 0.465 | 0.418 | 0.378 | 0.368 | 0.383 | 0.405 | 0.359 | 0.369 | 0.393 | 0.381 | 0.391 | 0.41 | 0.426 |
PVTv2 | 0.51 | 0.431 | 0.41 | 0.403 | 0.4 | 0.395 | 0.414 | 0.389 | 0.41 | 0.384 | 0.347 | 0.382 | 0.392 | 0.436 |
Clean |
Random |
AdvCam |
UPC |
NatPatch |
MTD |
LAP |
InvisCloak |
DAP |
AdvTshirt |
AdvTexture |
AdvPatch |
AdvPattern |
AdvCaT |
ATSS | 0.157 | 0.148 | 0.147 | 0.121 | 0.113 | 0.127 | 0.155 | 0.138 | 0.143 | 0.128 | 0.098 | 0.132 | 0.138 | 0.151 |
AutoAssign | 0.149 | 0.137 | 0.13 | 0.082 | 0.098 | 0.12 | 0.136 | 0.126 | 0.113 | 0.114 | 0.089 | 0.123 | 0.131 | 0.143 |
CenterNet | 0.161 | 0.145 | 0.147 | 0.12 | 0.111 | 0.141 | 0.151 | 0.132 | 0.132 | 0.135 | 0.102 | 0.124 | 0.142 | 0.166 |
CentripetalNet | 0.143 | 0.145 | 0.144 | 0.142 | 0.083 | 0.145 | 0.139 | 0.142 | 0.131 | 0.128 | 0.109 | 0.127 | 0.138 | 0.143 |
CornerNet | 0.142 | 0.141 | 0.138 | 0.12 | 0.07 | 0.133 | 0.141 | 0.128 | 0.125 | 0.115 | 0.091 | 0.111 | 0.132 | 0.14 |
DDOD | 0.142 | 0.141 | 0.129 | 0.102 | 0.119 | 0.115 | 0.139 | 0.133 | 0.131 | 0.129 | 0.085 | 0.114 | 0.126 | 0.128 |
DyHead | 0.135 | 0.137 | 0.144 | 0.113 | 0.109 | 0.127 | 0.149 | 0.132 | 0.141 | 0.12 | 0.106 | 0.122 | 0.131 | 0.134 |
EfficientNet | 0.123 | 0.115 | 0.119 | 0.11 | 0.113 | 0.106 | 0.116 | 0.105 | 0.11 | 0.106 | 0.104 | 0.104 | 0.109 | 0.115 |
FCOS | 0.126 | 0.117 | 0.117 | 0.1 | 0.102 | 0.108 | 0.118 | 0.109 | 0.106 | 0.107 | 0.094 | 0.106 | 0.109 | 0.123 |
FoveaBox | 0.173 | 0.163 | 0.166 | 0.135 | 0.133 | 0.141 | 0.162 | 0.141 | 0.155 | 0.136 | 0.098 | 0.131 | 0.15 | 0.161 |
FreeAnchor | 0.165 | 0.155 | 0.147 | 0.122 | 0.106 | 0.121 | 0.149 | 0.134 | 0.132 | 0.108 | 0.083 | 0.127 | 0.14 | 0.152 |
FSAF | 0.175 | 0.174 | 0.166 | 0.135 | 0.13 | 0.142 | 0.169 | 0.155 | 0.158 | 0.131 | 0.097 | 0.144 | 0.159 | 0.167 |
GFL | 0.177 | 0.164 | 0.162 | 0.123 | 0.13 | 0.138 | 0.16 | 0.143 | 0.157 | 0.144 | 0.108 | 0.133 | 0.153 | 0.164 |
LD | 0.179 | 0.164 | 0.162 | 0.114 | 0.109 | 0.143 | 0.162 | 0.147 | 0.148 | 0.142 | 0.101 | 0.145 | 0.157 | 0.162 |
NAS-FPN | 0.127 | 0.119 | 0.122 | 0.107 | 0.11 | 0.106 | 0.133 | 0.112 | 0.111 | 0.106 | 0.076 | 0.11 | 0.111 | 0.122 |
PAA | 0.132 | 0.132 | 0.13 | 0.112 | 0.103 | 0.116 | 0.135 | 0.128 | 0.129 | 0.108 | 0.081 | 0.12 | 0.122 | 0.129 |
RetinaNet | 0.155 | 0.167 | 0.151 | 0.135 | 0.123 | 0.13 | 0.158 | 0.152 | 0.142 | 0.134 | 0.106 | 0.144 | 0.144 | 0.154 |
RTMDet | 0.152 | 0.133 | 0.148 | 0.158 | 0.129 | 0.12 | 0.14 | 0.112 | 0.137 | 0.129 | 0.126 | 0.117 | 0.121 | 0.138 |
TOOD | 0.131 | 0.145 | 0.139 | 0.108 | 0.099 | 0.121 | 0.154 | 0.13 | 0.143 | 0.122 | 0.093 | 0.122 | 0.123 | 0.132 |
VarifocalNet | 0.141 | 0.152 | 0.145 | 0.11 | 0.111 | 0.127 | 0.157 | 0.133 | 0.144 | 0.125 | 0.095 | 0.132 | 0.137 | 0.154 |
YOLOv5 | 0.135 | 0.124 | 0.132 | 0.129 | 0.104 | 0.122 | 0.136 | 0.124 | 0.142 | 0.108 | 0.082 | 0.106 | 0.122 | 0.124 |
YOLOv6 | 0.13 | 0.121 | 0.129 | 0.139 | 0.119 | 0.118 | 0.126 | 0.123 | 0.128 | 0.118 | 0.115 | 0.116 | 0.119 | 0.123 |
YOLOv7 | 0.125 | 0.115 | 0.129 | 0.117 | 0.083 | 0.101 | 0.115 | 0.105 | 0.129 | 0.101 | 0.091 | 0.091 | 0.111 | 0.115 |
YOLOv8 | 0.12 | 0.113 | 0.12 | 0.128 | 0.111 | 0.111 | 0.115 | 0.111 | 0.119 | 0.109 | 0.106 | 0.106 | 0.111 | 0.112 |
YOLOX | 0.13 | 0.122 | 0.129 | 0.138 | 0.106 | 0.117 | 0.129 | 0.12 | 0.137 | 0.121 | 0.103 | 0.112 | 0.117 | 0.123 |
Faster R-CNN | 0.159 | 0.166 | 0.146 | 0.114 | 0.11 | 0.119 | 0.164 | 0.136 | 0.133 | 0.122 | 0.083 | 0.121 | 0.152 | 0.158 |
Cascade R-CNN | 0.171 | 0.166 | 0.165 | 0.12 | 0.122 | 0.141 | 0.17 | 0.136 | 0.15 | 0.136 | 0.09 | 0.131 | 0.154 | 0.165 |
Cascade RPN | 0.163 | 0.161 | 0.153 | 0.105 | 0.109 | 0.138 | 0.153 | 0.14 | 0.142 | 0.13 | 0.082 | 0.129 | 0.151 | 0.162 |
Double Heads | 0.169 | 0.161 | 0.164 | 0.12 | 0.11 | 0.134 | 0.165 | 0.128 | 0.144 | 0.128 | 0.085 | 0.126 | 0.151 | 0.16 |
FPG | 0.133 | 0.136 | 0.132 | 0.121 | 0.109 | 0.112 | 0.141 | 0.115 | 0.124 | 0.105 | 0.073 | 0.112 | 0.123 | 0.126 |
Grid R-CNN | 0.151 | 0.145 | 0.145 | 0.109 | 0.11 | 0.123 | 0.15 | 0.13 | 0.139 | 0.116 | 0.091 | 0.114 | 0.137 | 0.144 |
Guided Anchoring | 0.142 | 0.162 | 0.153 | 0.119 | 0.107 | 0.137 | 0.163 | 0.136 | 0.132 | 0.123 | 0.094 | 0.128 | 0.146 | 0.151 |
HRNet | 0.147 | 0.139 | 0.146 | 0.132 | 0.105 | 0.117 | 0.141 | 0.133 | 0.122 | 0.132 | 0.109 | 0.117 | 0.122 | 0.144 |
Libra R-CNN | 0.155 | 0.145 | 0.136 | 0.133 | 0.119 | 0.119 | 0.134 | 0.123 | 0.126 | 0.121 | 0.102 | 0.115 | 0.12 | 0.135 |
PAFPN | 0.163 | 0.157 | 0.158 | 0.118 | 0.12 | 0.124 | 0.156 | 0.137 | 0.137 | 0.125 | 0.086 | 0.121 | 0.149 | 0.165 |
RepPoints | 0.178 | 0.177 | 0.166 | 0.124 | 0.14 | 0.131 | 0.167 | 0.135 | 0.146 | 0.132 | 0.099 | 0.129 | 0.154 | 0.174 |
Res2Net | 0.123 | 0.115 | 0.122 | 0.112 | 0.075 | 0.106 | 0.132 | 0.113 | 0.115 | 0.1 | 0.086 | 0.099 | 0.111 | 0.12 |
ResNeSt | 0.125 | 0.132 | 0.118 | 0.11 | 0.11 | 0.109 | 0.125 | 0.108 | 0.102 | 0.101 | 0.091 | 0.1 | 0.118 | 0.127 |
SABL | 0.177 | 0.169 | 0.166 | 0.118 | 0.133 | 0.145 | 0.166 | 0.151 | 0.147 | 0.143 | 0.096 | 0.141 | 0.161 | 0.172 |
Sparse R-CNN | 0.135 | 0.134 | 0.138 | 0.104 | 0.092 | 0.114 | 0.141 | 0.106 | 0.111 | 0.109 | 0.095 | 0.103 | 0.122 | 0.131 |
DETR | 0.171 | 0.151 | 0.151 | 0.101 | 0.088 | 0.142 | 0.152 | 0.146 | 0.126 | 0.138 | 0.095 | 0.128 | 0.142 | 0.151 |
Conditional DETR | 0.157 | 0.138 | 0.128 | 0.1 | 0.09 | 0.114 | 0.129 | 0.112 | 0.123 | 0.108 | 0.068 | 0.105 | 0.111 | 0.135 |
DDQ | 0.119 | 0.119 | 0.124 | 0.099 | 0.099 | 0.112 | 0.124 | 0.116 | 0.113 | 0.11 | 0.085 | 0.113 | 0.117 | 0.12 |
DAB-DETR | 0.118 | 0.111 | 0.117 | 0.092 | 0.094 | 0.099 | 0.111 | 0.095 | 0.101 | 0.098 | 0.071 | 0.087 | 0.103 | 0.117 |
Deformable DETR | 0.136 | 0.129 | 0.142 | 0.093 | 0.107 | 0.118 | 0.134 | 0.128 | 0.12 | 0.12 | 0.073 | 0.114 | 0.137 | 0.121 |
DINO | 0.112 | 0.11 | 0.112 | 0.086 | 0.068 | 0.097 | 0.113 | 0.106 | 0.098 | 0.101 | 0.083 | 0.097 | 0.1 | 0.112 |
PVT | 0.14 | 0.134 | 0.117 | 0.107 | 0.099 | 0.108 | 0.115 | 0.099 | 0.102 | 0.111 | 0.106 | 0.109 | 0.113 | 0.12 |
PVTv2 | 0.156 | 0.132 | 0.119 | 0.119 | 0.112 | 0.108 | 0.126 | 0.118 | 0.122 | 0.11 | 0.097 | 0.108 | 0.111 | 0.129 |
Clean |
Random |
AdvCam |
UPC |
NatPatch |
MTD |
LAP |
InvisCloak |
DAP |
AdvTshirt |
AdvTexture |
AdvPatch |
AdvPattern |
AdvCaT |
ATSS | 0.835 | 0.827 | 0.823 | 0.802 | 0.782 | 0.823 | 0.818 | 0.789 | 0.788 | 0.798 | 0.741 | 0.786 | 0.823 | 0.844 |
AutoAssign | 0.854 | 0.837 | 0.841 | 0.794 | 0.827 | 0.832 | 0.826 | 0.814 | 0.817 | 0.827 | 0.793 | 0.811 | 0.832 | 0.851 |
CenterNet | 0.848 | 0.835 | 0.84 | 0.849 | 0.809 | 0.842 | 0.829 | 0.823 | 0.822 | 0.828 | 0.794 | 0.82 | 0.848 | 0.858 |
CentripetalNet | 0.854 | 0.858 | 0.838 | 0.809 | 0.777 | 0.85 | 0.83 | 0.86 | 0.82 | 0.814 | 0.737 | 0.841 | 0.852 | 0.854 |
CornerNet | 0.871 | 0.87 | 0.853 | 0.814 | 0.778 | 0.853 | 0.841 | 0.848 | 0.826 | 0.805 | 0.74 | 0.817 | 0.857 | 0.858 |
DDOD | 0.737 | 0.729 | 0.739 | 0.723 | 0.746 | 0.716 | 0.732 | 0.713 | 0.725 | 0.72 | 0.678 | 0.694 | 0.72 | 0.742 |
DyHead | 0.725 | 0.731 | 0.747 | 0.702 | 0.7 | 0.723 | 0.745 | 0.711 | 0.711 | 0.725 | 0.724 | 0.716 | 0.727 | 0.744 |
EfficientNet | 0.794 | 0.771 | 0.789 | 0.768 | 0.751 | 0.758 | 0.781 | 0.762 | 0.754 | 0.762 | 0.748 | 0.752 | 0.766 | 0.793 |
FCOS | 0.897 | 0.905 | 0.89 | 0.864 | 0.86 | 0.891 | 0.894 | 0.893 | 0.858 | 0.897 | 0.82 | 0.898 | 0.899 | 0.908 |
FoveaBox | 0.824 | 0.814 | 0.836 | 0.836 | 0.794 | 0.825 | 0.813 | 0.799 | 0.819 | 0.81 | 0.785 | 0.776 | 0.818 | 0.836 |
FreeAnchor | 0.792 | 0.801 | 0.788 | 0.791 | 0.772 | 0.786 | 0.789 | 0.768 | 0.769 | 0.781 | 0.717 | 0.784 | 0.793 | 0.809 |
FSAF | 0.812 | 0.814 | 0.818 | 0.803 | 0.79 | 0.807 | 0.822 | 0.808 | 0.808 | 0.803 | 0.785 | 0.803 | 0.805 | 0.812 |
GFL | 0.827 | 0.803 | 0.805 | 0.796 | 0.782 | 0.806 | 0.784 | 0.781 | 0.763 | 0.785 | 0.743 | 0.777 | 0.804 | 0.823 |
LD | 0.809 | 0.797 | 0.802 | 0.785 | 0.762 | 0.787 | 0.786 | 0.764 | 0.77 | 0.777 | 0.735 | 0.763 | 0.797 | 0.813 |
NAS-FPN | 0.773 | 0.765 | 0.775 | 0.726 | 0.75 | 0.74 | 0.768 | 0.744 | 0.732 | 0.741 | 0.7 | 0.751 | 0.76 | 0.778 |
PAA | 0.816 | 0.819 | 0.821 | 0.808 | 0.791 | 0.805 | 0.805 | 0.792 | 0.784 | 0.78 | 0.73 | 0.783 | 0.807 | 0.837 |
RetinaNet | 0.838 | 0.823 | 0.843 | 0.817 | 0.795 | 0.823 | 0.822 | 0.802 | 0.796 | 0.818 | 0.777 | 0.791 | 0.823 | 0.848 |
RTMDet | 0.976 | 0.978 | 0.974 | 0.973 | 0.972 | 0.975 | 0.974 | 0.966 | 0.971 | 0.97 | 0.966 | 0.976 | 0.977 | 0.976 |
TOOD | 0.73 | 0.732 | 0.731 | 0.708 | 0.691 | 0.734 | 0.737 | 0.716 | 0.723 | 0.724 | 0.69 | 0.714 | 0.723 | 0.733 |
VarifocalNet | 0.794 | 0.772 | 0.786 | 0.767 | 0.738 | 0.765 | 0.77 | 0.744 | 0.755 | 0.758 | 0.729 | 0.745 | 0.771 | 0.791 |
YOLOv5 | 0.814 | 0.832 | 0.836 | 0.773 | 0.782 | 0.825 | 0.831 | 0.841 | 0.814 | 0.829 | 0.817 | 0.836 | 0.831 | 0.843 |
YOLOv6 | 0.96 | 0.963 | 0.957 | 0.951 | 0.939 | 0.959 | 0.943 | 0.964 | 0.943 | 0.95 | 0.921 | 0.956 | 0.963 | 0.966 |
YOLOv7 | 0.729 | 0.73 | 0.744 | 0.723 | 0.7 | 0.716 | 0.731 | 0.712 | 0.721 | 0.731 | 0.721 | 0.723 | 0.722 | 0.743 |
YOLOv8 | 0.677 | 0.667 | 0.686 | 0.67 | 0.669 | 0.676 | 0.676 | 0.679 | 0.687 | 0.674 | 0.675 | 0.674 | 0.668 | 0.679 |
YOLOX | 0.685 | 0.697 | 0.702 | 0.702 | 0.702 | 0.7 | 0.708 | 0.706 | 0.709 | 0.707 | 0.731 | 0.71 | 0.682 | 0.692 |
Faster R-CNN | 0.69 | 0.682 | 0.691 | 0.625 | 0.659 | 0.679 | 0.674 | 0.653 | 0.645 | 0.679 | 0.615 | 0.639 | 0.67 | 0.685 |
Cascade R-CNN | 0.699 | 0.689 | 0.703 | 0.644 | 0.652 | 0.679 | 0.695 | 0.66 | 0.673 | 0.677 | 0.623 | 0.651 | 0.68 | 0.7 |
Cascade RPN | 0.785 | 0.791 | 0.789 | 0.759 | 0.758 | 0.78 | 0.776 | 0.748 | 0.751 | 0.774 | 0.715 | 0.758 | 0.784 | 0.797 |
Double Heads | 0.696 | 0.688 | 0.7 | 0.636 | 0.639 | 0.677 | 0.678 | 0.664 | 0.654 | 0.679 | 0.623 | 0.643 | 0.676 | 0.696 |
FPG | 0.656 | 0.659 | 0.665 | 0.598 | 0.598 | 0.643 | 0.659 | 0.636 | 0.622 | 0.642 | 0.588 | 0.635 | 0.659 | 0.661 |
Grid R-CNN | 0.677 | 0.664 | 0.67 | 0.638 | 0.635 | 0.666 | 0.669 | 0.654 | 0.655 | 0.659 | 0.632 | 0.644 | 0.657 | 0.661 |
Guided Anchoring | 0.74 | 0.732 | 0.748 | 0.729 | 0.709 | 0.738 | 0.745 | 0.733 | 0.711 | 0.736 | 0.72 | 0.718 | 0.739 | 0.739 |
HRNet | 0.665 | 0.669 | 0.679 | 0.645 | 0.644 | 0.671 | 0.673 | 0.661 | 0.646 | 0.676 | 0.65 | 0.645 | 0.649 | 0.678 |
Libra R-CNN | 0.829 | 0.813 | 0.831 | 0.822 | 0.766 | 0.81 | 0.811 | 0.781 | 0.776 | 0.797 | 0.754 | 0.778 | 0.819 | 0.848 |
PAFPN | 0.685 | 0.682 | 0.697 | 0.644 | 0.647 | 0.678 | 0.679 | 0.664 | 0.658 | 0.682 | 0.614 | 0.651 | 0.679 | 0.698 |
RepPoints | 0.826 | 0.813 | 0.822 | 0.816 | 0.813 | 0.804 | 0.807 | 0.799 | 0.808 | 0.801 | 0.783 | 0.779 | 0.813 | 0.828 |
Res2Net | 0.636 | 0.63 | 0.638 | 0.599 | 0.543 | 0.627 | 0.623 | 0.614 | 0.602 | 0.623 | 0.578 | 0.618 | 0.63 | 0.634 |
ResNeSt | 0.648 | 0.651 | 0.625 | 0.61 | 0.61 | 0.636 | 0.619 | 0.614 | 0.586 | 0.607 | 0.593 | 0.609 | 0.64 | 0.654 |
SABL | 0.711 | 0.704 | 0.707 | 0.65 | 0.66 | 0.69 | 0.678 | 0.675 | 0.675 | 0.688 | 0.635 | 0.661 | 0.696 | 0.706 |
Sparse R-CNN | 0.702 | 0.685 | 0.694 | 0.67 | 0.668 | 0.672 | 0.688 | 0.65 | 0.644 | 0.667 | 0.649 | 0.651 | 0.68 | 0.694 |
DETR | 0.893 | 0.9 | 0.877 | 0.761 | 0.798 | 0.904 | 0.887 | 0.891 | 0.827 | 0.882 | 0.789 | 0.889 | 0.895 | 0.892 |
Conditional DETR | 0.73 | 0.729 | 0.717 | 0.706 | 0.708 | 0.697 | 0.703 | 0.67 | 0.691 | 0.683 | 0.643 | 0.692 | 0.69 | 0.712 |
DDQ | 0.774 | 0.774 | 0.768 | 0.768 | 0.758 | 0.746 | 0.757 | 0.728 | 0.744 | 0.748 | 0.725 | 0.745 | 0.746 | 0.773 |
DAB-DETR | 0.743 | 0.739 | 0.745 | 0.737 | 0.76 | 0.729 | 0.742 | 0.705 | 0.73 | 0.717 | 0.703 | 0.716 | 0.722 | 0.743 |
Deformable DETR | 0.703 | 0.699 | 0.707 | 0.68 | 0.693 | 0.671 | 0.692 | 0.666 | 0.675 | 0.673 | 0.591 | 0.67 | 0.679 | 0.693 |
DINO | 0.739 | 0.748 | 0.744 | 0.735 | 0.747 | 0.732 | 0.747 | 0.723 | 0.72 | 0.727 | 0.718 | 0.734 | 0.73 | 0.736 |
PVT | 0.703 | 0.705 | 0.7 | 0.692 | 0.673 | 0.693 | 0.696 | 0.679 | 0.669 | 0.692 | 0.681 | 0.691 | 0.688 | 0.699 |
PVTv2 | 0.738 | 0.733 | 0.714 | 0.736 | 0.731 | 0.712 | 0.733 | 0.715 | 0.705 | 0.725 | 0.7 | 0.716 | 0.709 | 0.731 |
Clean |
Random |
AdvCam |
UPC |
NatPatch |
MTD |
LAP |
InvisCloak |
DAP |
AdvTshirt |
AdvTexture |
AdvPatch |
AdvPattern |
AdvCaT |
ATSS | 0.326 | 0.313 | 0.318 | 0.301 | 0.298 | 0.304 | 0.306 | 0.291 | 0.302 | 0.297 | 0.269 | 0.286 | 0.307 | 0.322 |
AutoAssign | 0.317 | 0.309 | 0.309 | 0.288 | 0.306 | 0.299 | 0.305 | 0.297 | 0.3 | 0.301 | 0.282 | 0.297 | 0.306 | 0.314 |
CenterNet | 0.351 | 0.344 | 0.342 | 0.341 | 0.315 | 0.342 | 0.332 | 0.326 | 0.325 | 0.332 | 0.309 | 0.329 | 0.346 | 0.352 |
CentripetalNet | 0.332 | 0.337 | 0.326 | 0.32 | 0.293 | 0.333 | 0.319 | 0.338 | 0.311 | 0.307 | 0.277 | 0.321 | 0.331 | 0.332 |
CornerNet | 0.331 | 0.333 | 0.325 | 0.306 | 0.281 | 0.322 | 0.314 | 0.314 | 0.306 | 0.297 | 0.27 | 0.298 | 0.321 | 0.326 |
DDOD | 0.273 | 0.267 | 0.275 | 0.262 | 0.28 | 0.256 | 0.267 | 0.256 | 0.264 | 0.262 | 0.245 | 0.251 | 0.262 | 0.273 |
DyHead | 0.283 | 0.284 | 0.294 | 0.267 | 0.273 | 0.275 | 0.293 | 0.273 | 0.279 | 0.279 | 0.27 | 0.274 | 0.279 | 0.291 |
EfficientNet | 0.29 | 0.283 | 0.295 | 0.283 | 0.284 | 0.277 | 0.291 | 0.278 | 0.284 | 0.28 | 0.274 | 0.272 | 0.276 | 0.291 |
FCOS | 0.339 | 0.341 | 0.341 | 0.323 | 0.331 | 0.331 | 0.339 | 0.337 | 0.322 | 0.337 | 0.296 | 0.335 | 0.333 | 0.344 |
FoveaBox | 0.304 | 0.298 | 0.316 | 0.309 | 0.29 | 0.295 | 0.293 | 0.285 | 0.301 | 0.292 | 0.276 | 0.279 | 0.294 | 0.309 |
FreeAnchor | 0.289 | 0.289 | 0.29 | 0.283 | 0.278 | 0.28 | 0.284 | 0.276 | 0.278 | 0.277 | 0.249 | 0.28 | 0.284 | 0.292 |
FSAF | 0.307 | 0.305 | 0.31 | 0.298 | 0.304 | 0.299 | 0.307 | 0.299 | 0.307 | 0.294 | 0.279 | 0.297 | 0.301 | 0.305 |
GFL | 0.311 | 0.298 | 0.304 | 0.29 | 0.29 | 0.292 | 0.29 | 0.283 | 0.285 | 0.284 | 0.267 | 0.281 | 0.295 | 0.308 |
LD | 0.302 | 0.293 | 0.298 | 0.283 | 0.28 | 0.285 | 0.291 | 0.278 | 0.286 | 0.283 | 0.263 | 0.276 | 0.29 | 0.297 |
NAS-FPN | 0.294 | 0.28 | 0.293 | 0.265 | 0.278 | 0.27 | 0.285 | 0.271 | 0.272 | 0.271 | 0.242 | 0.272 | 0.278 | 0.288 |
PAA | 0.321 | 0.322 | 0.323 | 0.308 | 0.303 | 0.312 | 0.32 | 0.312 | 0.313 | 0.303 | 0.269 | 0.304 | 0.318 | 0.33 |
RetinaNet | 0.317 | 0.32 | 0.327 | 0.314 | 0.305 | 0.312 | 0.316 | 0.309 | 0.307 | 0.312 | 0.295 | 0.304 | 0.312 | 0.327 |
RTMDet | 0.246 | 0.242 | 0.248 | 0.235 | 0.233 | 0.236 | 0.244 | 0.229 | 0.237 | 0.238 | 0.233 | 0.233 | 0.234 | 0.238 |
TOOD | 0.277 | 0.276 | 0.278 | 0.262 | 0.258 | 0.271 | 0.277 | 0.266 | 0.274 | 0.269 | 0.249 | 0.265 | 0.268 | 0.278 |
VarifocalNet | 0.295 | 0.287 | 0.296 | 0.281 | 0.277 | 0.28 | 0.286 | 0.271 | 0.28 | 0.277 | 0.262 | 0.274 | 0.284 | 0.294 |
YOLOv5 | 0.241 | 0.239 | 0.243 | 0.235 | 0.222 | 0.237 | 0.242 | 0.238 | 0.241 | 0.229 | 0.179 | 0.223 | 0.238 | 0.235 |
YOLOv6 | 0.241 | 0.237 | 0.24 | 0.232 | 0.233 | 0.231 | 0.236 | 0.232 | 0.238 | 0.231 | 0.228 | 0.228 | 0.232 | 0.233 |
YOLOv7 | 0.242 | 0.236 | 0.242 | 0.229 | 0.207 | 0.221 | 0.237 | 0.218 | 0.235 | 0.226 | 0.215 | 0.207 | 0.232 | 0.233 |
YOLOv8 | 0.244 | 0.241 | 0.244 | 0.235 | 0.235 | 0.237 | 0.241 | 0.238 | 0.242 | 0.237 | 0.232 | 0.233 | 0.239 | 0.238 |
YOLOX | 0.242 | 0.241 | 0.243 | 0.233 | 0.217 | 0.237 | 0.241 | 0.237 | 0.239 | 0.236 | 0.227 | 0.234 | 0.239 | 0.237 |
Faster R-CNN | 0.256 | 0.255 | 0.261 | 0.228 | 0.242 | 0.246 | 0.253 | 0.24 | 0.244 | 0.25 | 0.212 | 0.233 | 0.247 | 0.255 |
Cascade R-CNN | 0.262 | 0.255 | 0.267 | 0.238 | 0.246 | 0.248 | 0.261 | 0.244 | 0.256 | 0.249 | 0.224 | 0.24 | 0.249 | 0.258 |
Cascade RPN | 0.279 | 0.278 | 0.28 | 0.26 | 0.266 | 0.27 | 0.274 | 0.261 | 0.266 | 0.268 | 0.241 | 0.262 | 0.273 | 0.281 |
Double Heads | 0.26 | 0.255 | 0.262 | 0.233 | 0.232 | 0.243 | 0.253 | 0.24 | 0.245 | 0.244 | 0.209 | 0.233 | 0.249 | 0.258 |
FPG | 0.248 | 0.243 | 0.251 | 0.217 | 0.211 | 0.233 | 0.247 | 0.235 | 0.235 | 0.231 | 0.196 | 0.229 | 0.242 | 0.241 |
Grid R-CNN | 0.25 | 0.244 | 0.251 | 0.235 | 0.24 | 0.243 | 0.25 | 0.241 | 0.248 | 0.246 | 0.234 | 0.238 | 0.24 | 0.244 |
Guided Anchoring | 0.271 | 0.266 | 0.271 | 0.254 | 0.252 | 0.262 | 0.272 | 0.259 | 0.257 | 0.265 | 0.244 | 0.253 | 0.264 | 0.268 |
HRNet | 0.253 | 0.248 | 0.256 | 0.243 | 0.233 | 0.246 | 0.25 | 0.246 | 0.243 | 0.252 | 0.239 | 0.234 | 0.238 | 0.254 |
Libra R-CNN | 0.321 | 0.314 | 0.319 | 0.319 | 0.293 | 0.302 | 0.307 | 0.294 | 0.3 | 0.303 | 0.285 | 0.291 | 0.309 | 0.326 |
PAFPN | 0.257 | 0.253 | 0.263 | 0.232 | 0.243 | 0.245 | 0.255 | 0.243 | 0.251 | 0.249 | 0.212 | 0.235 | 0.25 | 0.259 |
RepPoints | 0.314 | 0.31 | 0.319 | 0.304 | 0.306 | 0.295 | 0.301 | 0.291 | 0.305 | 0.298 | 0.28 | 0.287 | 0.306 | 0.316 |
Res2Net | 0.237 | 0.231 | 0.24 | 0.22 | 0.193 | 0.229 | 0.234 | 0.227 | 0.226 | 0.228 | 0.216 | 0.222 | 0.23 | 0.232 |
ResNeSt | 0.244 | 0.248 | 0.236 | 0.224 | 0.226 | 0.236 | 0.236 | 0.231 | 0.217 | 0.226 | 0.214 | 0.221 | 0.241 | 0.245 |
SABL | 0.271 | 0.26 | 0.27 | 0.24 | 0.251 | 0.254 | 0.256 | 0.247 | 0.257 | 0.252 | 0.228 | 0.245 | 0.26 | 0.265 |
Sparse R-CNN | 0.268 | 0.262 | 0.267 | 0.25 | 0.253 | 0.251 | 0.263 | 0.246 | 0.247 | 0.253 | 0.247 | 0.244 | 0.259 | 0.265 |
DETR | 0.42 | 0.42 | 0.391 | 0.334 | 0.352 | 0.399 | 0.39 | 0.399 | 0.351 | 0.383 | 0.328 | 0.392 | 0.404 | 0.397 |
Conditional DETR | 0.293 | 0.302 | 0.29 | 0.278 | 0.281 | 0.285 | 0.28 | 0.268 | 0.269 | 0.268 | 0.236 | 0.275 | 0.278 | 0.286 |
DDQ | 0.322 | 0.32 | 0.32 | 0.303 | 0.304 | 0.301 | 0.313 | 0.29 | 0.29 | 0.305 | 0.284 | 0.299 | 0.3 | 0.322 |
DAB-DETR | 0.301 | 0.301 | 0.303 | 0.3 | 0.31 | 0.295 | 0.296 | 0.277 | 0.294 | 0.285 | 0.27 | 0.28 | 0.292 | 0.303 |
Deformable DETR | 0.281 | 0.271 | 0.278 | 0.258 | 0.282 | 0.256 | 0.27 | 0.252 | 0.261 | 0.26 | 0.221 | 0.257 | 0.261 | 0.267 |
DINO | 0.303 | 0.309 | 0.303 | 0.29 | 0.294 | 0.294 | 0.302 | 0.287 | 0.276 | 0.293 | 0.276 | 0.293 | 0.295 | 0.299 |
PVT | 0.273 | 0.275 | 0.27 | 0.264 | 0.253 | 0.268 | 0.271 | 0.26 | 0.254 | 0.271 | 0.266 | 0.267 | 0.266 | 0.275 |
PVTv2 | 0.284 | 0.283 | 0.277 | 0.278 | 0.277 | 0.271 | 0.281 | 0.276 | 0.266 | 0.279 | 0.264 | 0.273 | 0.273 | 0.28 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.975 | 0.963 | 0.863 | 0.963 | 0.95 | 1.0 | 0.863 | 0.963 | 0.887 | 0.875 |
AutoAssign | 0.975 | 1.0 | 0.975 | 1.0 | 1.0 | 1.0 | 0.975 | 0.988 | 1.0 | 0.988 |
CenterNet | 0.95 | 1.0 | 0.963 | 1.0 | 0.975 | 1.0 | 0.887 | 1.0 | 1.0 | 0.925 |
CentripetalNet | 0.975 | 1.0 | 0.938 | 1.0 | 0.95 | 1.0 | 0.975 | 1.0 | 0.988 | 0.988 |
CornerNet | 1.0 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.95 |
DDOD | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 |
DyHead | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 0.975 | 1.0 |
EfficientNet | 0.938 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 | 0.975 |
FCOS | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 0.975 |
FoveaBox | 1.0 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 |
FreeAnchor | 1.0 | 0.863 | 0.812 | 0.975 | 0.863 | 1.0 | 0.8 | 0.875 | 0.875 | 0.887 |
FSAF | 0.988 | 0.887 | 0.887 | 0.988 | 0.925 | 1.0 | 0.925 | 0.975 | 0.9 | 0.912 |
GFL | 0.988 | 0.875 | 0.875 | 1.0 | 0.912 | 1.0 | 0.887 | 0.875 | 0.875 | 0.875 |
LD | 1.0 | 0.9 | 0.887 | 1.0 | 0.975 | 1.0 | 0.887 | 1.0 | 0.9 | 0.975 |
NAS-FPN | 1.0 | 0.912 | 1.0 | 1.0 | 0.963 | 1.0 | 0.975 | 1.0 | 0.975 | 0.963 |
PAA | 0.988 | 1.0 | 0.988 | 1.0 | 0.975 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 |
RetinaNet | 0.988 | 0.938 | 0.925 | 0.963 | 0.925 | 1.0 | 0.863 | 0.988 | 0.863 | 0.925 |
RTMDet | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 |
TOOD | 1.0 | 1.0 | 0.912 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 |
VarifocalNet | 0.988 | 0.938 | 0.85 | 0.887 | 0.95 | 1.0 | 0.925 | 0.85 | 0.875 | 0.887 |
YOLOv5 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv6 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.963 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv7 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.925 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv8 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.975 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOX | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 |
Faster R-CNN | 0.988 | 0.85 | 0.562 | 0.775 | 0.85 | 0.938 | 0.762 | 0.887 | 0.725 | 0.875 |
Cascade R-CNN | 1.0 | 0.863 | 0.738 | 0.875 | 0.912 | 0.975 | 0.787 | 0.925 | 0.787 | 0.863 |
Cascade RPN | 1.0 | 1.0 | 0.988 | 1.0 | 0.988 | 0.975 | 0.988 | 1.0 | 0.988 | 0.988 |
Double Heads | 0.975 | 0.9 | 0.812 | 0.875 | 0.825 | 1.0 | 0.838 | 0.975 | 0.838 | 0.875 |
FPG | 0.988 | 0.988 | 0.95 | 1.0 | 0.988 | 1.0 | 0.9 | 0.988 | 0.938 | 0.925 |
Grid R-CNN | 1.0 | 0.863 | 0.738 | 0.963 | 0.85 | 0.988 | 0.838 | 1.0 | 0.875 | 0.875 |
Guided Anchoring | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 | 0.975 |
HRNet | 0.938 | 0.875 | 0.938 | 0.95 | 0.875 | 0.963 | 0.775 | 0.9 | 0.95 | 0.925 |
Libra R-CNN | 0.988 | 0.938 | 0.938 | 0.963 | 0.963 | 1.0 | 0.975 | 1.0 | 0.9 | 0.975 |
PAFPN | 0.988 | 0.863 | 0.637 | 0.825 | 0.887 | 0.963 | 0.838 | 0.963 | 0.825 | 0.863 |
RepPoints | 0.988 | 0.975 | 0.875 | 1.0 | 0.9 | 1.0 | 0.975 | 0.975 | 0.975 | 0.912 |
Res2Net | 0.975 | 1.0 | 0.925 | 0.988 | 0.988 | 1.0 | 0.912 | 0.938 | 0.988 | 1.0 |
ResNeSt | 0.975 | 0.863 | 0.825 | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 | 0.863 | 0.963 |
SABL | 1.0 | 0.875 | 0.775 | 0.938 | 0.887 | 0.988 | 0.787 | 0.95 | 0.85 | 0.863 |
Sparse R-CNN | 0.975 | 0.912 | 0.85 | 1.0 | 0.988 | 0.988 | 0.95 | 0.95 | 0.838 | 0.9 |
DETR | 0.95 | 0.537 | 0.388 | 0.237 | 0.8 | 0.713 | 0.812 | 0.5 | 0.487 | 0.8 |
Conditional DETR | 0.988 | 0.887 | 0.975 | 0.975 | 0.975 | 1.0 | 1.0 | 1.0 | 0.875 | 0.912 |
DDQ | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.975 | 1.0 |
Deformable DETR | 0.963 | 0.975 | 0.9 | 0.938 | 0.95 | 0.975 | 0.9 | 0.975 | 0.912 | 0.938 |
DINO | 1.0 | 1.0 | 0.988 | 0.925 | 0.975 | 0.988 | 1.0 | 1.0 | 0.9 | 1.0 |
PVT | 0.988 | 1.0 | 0.75 | 0.988 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 0.925 |
PVTv2 | 1.0 | 0.988 | 0.988 | 1.0 | 1.0 | 1.0 | 0.887 | 1.0 | 1.0 | 0.912 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 1.0 | 0.979 | 0.885 | 0.99 | 0.969 | 0.99 | 0.844 | 0.969 | 0.875 | 0.875 |
AutoAssign | 0.958 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 |
CenterNet | 0.99 | 0.979 | 0.99 | 0.99 | 0.979 | 0.99 | 0.885 | 0.979 | 0.99 | 0.938 |
CentripetalNet | 1.0 | 0.99 | 0.979 | 0.99 | 0.99 | 0.99 | 0.979 | 1.0 | 0.99 | 0.979 |
CornerNet | 0.99 | 0.99 | 0.958 | 0.99 | 0.99 | 0.99 | 0.979 | 1.0 | 0.969 | 0.99 |
DDOD | 1.0 | 1.0 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
DyHead | 1.0 | 1.0 | 0.979 | 1.0 | 0.99 | 1.0 | 1.0 | 0.979 | 0.969 | 1.0 |
EfficientNet | 0.958 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 | 0.979 | 1.0 | 0.99 | 0.979 |
FCOS | 1.0 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.958 |
FoveaBox | 0.99 | 0.99 | 0.875 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 |
FreeAnchor | 0.99 | 0.865 | 0.833 | 0.979 | 0.885 | 0.99 | 0.792 | 0.938 | 0.885 | 0.865 |
FSAF | 1.0 | 0.906 | 0.812 | 0.99 | 0.938 | 0.99 | 0.865 | 0.938 | 0.875 | 0.885 |
GFL | 0.99 | 0.927 | 0.969 | 0.99 | 0.938 | 0.99 | 0.885 | 0.927 | 0.917 | 0.896 |
LD | 0.99 | 0.948 | 0.896 | 0.99 | 0.958 | 0.99 | 0.823 | 0.979 | 0.896 | 0.99 |
NAS-FPN | 1.0 | 0.885 | 0.969 | 0.979 | 0.948 | 1.0 | 0.979 | 1.0 | 0.927 | 0.917 |
PAA | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.99 | 0.958 | 0.99 | 0.99 | 0.99 |
RetinaNet | 0.99 | 0.865 | 0.865 | 0.917 | 0.948 | 0.979 | 0.854 | 0.917 | 0.802 | 0.875 |
RTMDet | 1.0 | 1.0 | 0.99 | 1.0 | 0.99 | 0.979 | 0.99 | 0.99 | 1.0 | 0.99 |
TOOD | 0.99 | 0.99 | 0.927 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
VarifocalNet | 1.0 | 0.958 | 0.802 | 0.865 | 0.938 | 0.99 | 0.885 | 0.823 | 0.833 | 0.875 |
YOLOv5 | 1.0 | 0.99 | 0.99 | 1.0 | 0.99 | 1.0 | 1.0 | 0.99 | 0.99 | 0.969 |
YOLOv6 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.979 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv7 | 1.0 | 1.0 | 1.0 | 0.99 | 0.99 | 0.917 | 1.0 | 1.0 | 0.99 | 0.99 |
YOLOv8 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.958 | 1.0 | 1.0 | 0.99 | 1.0 |
YOLOX | 1.0 | 0.99 | 1.0 | 0.99 | 0.99 | 0.958 | 0.99 | 1.0 | 0.99 | 0.99 |
Faster R-CNN | 0.99 | 0.865 | 0.74 | 0.969 | 0.885 | 0.99 | 0.792 | 0.948 | 0.833 | 0.885 |
Cascade R-CNN | 0.99 | 0.865 | 0.854 | 0.969 | 0.885 | 0.99 | 0.823 | 0.917 | 0.844 | 0.854 |
Cascade RPN | 0.99 | 0.958 | 0.938 | 0.99 | 0.938 | 0.969 | 0.906 | 0.99 | 0.99 | 0.958 |
Double Heads | 0.99 | 0.896 | 0.927 | 0.969 | 0.823 | 0.99 | 0.865 | 0.99 | 0.844 | 0.865 |
FPG | 1.0 | 0.948 | 0.896 | 0.979 | 0.958 | 0.99 | 0.833 | 0.99 | 0.969 | 0.906 |
Grid R-CNN | 0.979 | 0.906 | 0.885 | 0.979 | 0.917 | 0.979 | 0.875 | 0.99 | 0.917 | 0.896 |
Guided Anchoring | 1.0 | 0.979 | 0.99 | 0.99 | 0.99 | 0.917 | 0.979 | 0.99 | 0.99 | 0.979 |
HRNet | 0.969 | 0.906 | 0.958 | 0.958 | 0.854 | 0.969 | 0.812 | 0.958 | 0.969 | 0.927 |
Libra R-CNN | 1.0 | 0.958 | 0.917 | 0.99 | 0.948 | 0.99 | 0.948 | 0.99 | 0.854 | 0.948 |
PAFPN | 0.99 | 0.875 | 0.74 | 0.938 | 0.938 | 0.969 | 0.927 | 0.979 | 0.844 | 0.844 |
RepPoints | 0.99 | 0.979 | 0.823 | 0.99 | 0.885 | 0.99 | 0.948 | 0.979 | 0.927 | 0.896 |
Res2Net | 0.979 | 0.979 | 0.979 | 0.99 | 0.99 | 0.99 | 0.927 | 0.99 | 0.979 | 0.979 |
ResNeSt | 0.958 | 0.865 | 0.938 | 0.979 | 0.99 | 0.99 | 0.99 | 0.979 | 0.917 | 0.969 |
SABL | 0.99 | 0.917 | 0.854 | 0.99 | 0.875 | 0.99 | 0.833 | 0.99 | 0.865 | 0.865 |
Sparse R-CNN | 0.99 | 0.969 | 0.917 | 0.99 | 0.99 | 0.99 | 0.958 | 0.979 | 0.875 | 0.948 |
DETR | 0.99 | 0.677 | 0.448 | 0.302 | 0.885 | 0.875 | 0.854 | 0.74 | 0.594 | 0.844 |
Conditional DETR | 0.99 | 0.958 | 0.938 | 0.979 | 0.979 | 0.948 | 0.969 | 0.99 | 0.906 | 0.896 |
DDQ | 1.0 | 0.99 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
Deformable DETR | 1.0 | 0.948 | 0.875 | 0.917 | 0.917 | 0.969 | 0.917 | 0.979 | 0.979 | 0.938 |
DINO | 1.0 | 0.99 | 0.948 | 0.938 | 0.979 | 0.979 | 0.979 | 0.99 | 0.927 | 0.99 |
PVT | 0.938 | 0.979 | 0.625 | 0.979 | 0.969 | 0.99 | 0.927 | 0.99 | 0.99 | 0.812 |
PVTv2 | 1.0 | 0.979 | 0.99 | 0.99 | 0.979 | 0.99 | 0.802 | 0.99 | 0.99 | 0.875 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.979 | 0.979 | 0.958 | 0.99 | 0.99 | 0.99 | 0.948 | 0.979 | 0.979 | 0.958 |
AutoAssign | 0.969 | 0.99 | 0.99 | 0.979 | 0.979 | 0.99 | 0.958 | 0.99 | 0.979 | 0.99 |
CenterNet | 0.99 | 0.979 | 1.0 | 0.99 | 0.979 | 1.0 | 0.99 | 1.0 | 1.0 | 0.969 |
CentripetalNet | 0.979 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
CornerNet | 1.0 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 |
DDOD | 0.99 | 0.979 | 0.958 | 0.99 | 0.979 | 0.99 | 0.969 | 0.99 | 0.99 | 0.979 |
DyHead | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.979 | 0.99 |
EfficientNet | 0.969 | 0.99 | 0.99 | 0.979 | 0.979 | 0.917 | 0.99 | 0.99 | 0.99 | 0.979 |
FCOS | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.969 | 0.99 | 0.99 | 0.969 |
FoveaBox | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 |
FreeAnchor | 0.979 | 0.917 | 0.865 | 0.979 | 0.927 | 0.99 | 0.802 | 0.969 | 0.802 | 0.917 |
FSAF | 1.0 | 0.917 | 0.885 | 0.99 | 0.958 | 0.979 | 0.896 | 0.99 | 0.917 | 0.906 |
GFL | 0.979 | 0.969 | 0.948 | 0.99 | 0.969 | 0.979 | 0.958 | 0.99 | 0.896 | 0.969 |
LD | 0.979 | 0.969 | 0.969 | 0.99 | 0.979 | 0.99 | 0.958 | 0.99 | 0.969 | 0.979 |
NAS-FPN | 0.99 | 0.969 | 0.979 | 0.979 | 0.979 | 0.99 | 0.99 | 0.99 | 0.948 | 0.958 |
PAA | 0.99 | 1.0 | 0.979 | 0.99 | 0.99 | 0.979 | 0.969 | 0.99 | 0.99 | 0.99 |
RetinaNet | 0.979 | 0.917 | 0.917 | 0.969 | 0.979 | 0.979 | 0.875 | 0.979 | 0.875 | 0.927 |
RTMDet | 0.99 | 0.99 | 1.0 | 0.938 | 0.99 | 0.958 | 1.0 | 1.0 | 0.99 | 0.99 |
TOOD | 0.99 | 0.99 | 0.948 | 0.979 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 |
VarifocalNet | 0.979 | 0.938 | 0.854 | 0.958 | 0.969 | 0.979 | 0.906 | 0.979 | 0.833 | 0.896 |
YOLOv5 | 0.99 | 1.0 | 1.0 | 0.99 | 1.0 | 0.979 | 0.99 | 1.0 | 0.99 | 1.0 |
YOLOv6 | 0.99 | 0.99 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 |
YOLOv7 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 | 0.99 |
YOLOv8 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
YOLOX | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
Faster R-CNN | 0.979 | 0.885 | 0.76 | 0.99 | 0.958 | 0.979 | 0.802 | 0.958 | 0.875 | 0.896 |
Cascade R-CNN | 0.99 | 0.906 | 0.823 | 0.979 | 0.948 | 0.99 | 0.792 | 0.958 | 0.875 | 0.896 |
Cascade RPN | 0.99 | 0.979 | 0.958 | 0.979 | 0.979 | 0.99 | 0.927 | 0.99 | 0.979 | 0.958 |
Double Heads | 0.99 | 0.948 | 0.885 | 0.99 | 0.896 | 0.99 | 0.875 | 0.969 | 0.927 | 0.885 |
FPG | 0.979 | 0.99 | 0.979 | 0.979 | 0.979 | 0.979 | 0.906 | 0.99 | 0.958 | 0.958 |
Grid R-CNN | 0.969 | 0.958 | 0.917 | 0.979 | 0.99 | 0.979 | 0.885 | 0.99 | 0.979 | 0.896 |
Guided Anchoring | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.958 | 0.99 | 0.99 | 0.99 | 0.979 |
HRNet | 0.979 | 0.927 | 0.875 | 0.979 | 0.938 | 0.979 | 0.885 | 0.969 | 0.99 | 0.969 |
Libra R-CNN | 0.99 | 0.969 | 0.969 | 0.979 | 0.979 | 0.979 | 0.969 | 0.99 | 0.958 | 0.938 |
PAFPN | 0.979 | 0.875 | 0.781 | 0.969 | 0.938 | 0.979 | 0.823 | 0.958 | 0.885 | 0.885 |
RepPoints | 0.99 | 0.958 | 0.896 | 0.99 | 0.948 | 0.99 | 0.979 | 1.0 | 1.0 | 0.938 |
Res2Net | 0.979 | 0.99 | 0.979 | 0.99 | 0.979 | 0.979 | 0.979 | 0.99 | 0.99 | 0.979 |
ResNeSt | 0.979 | 0.927 | 0.958 | 0.969 | 0.969 | 0.979 | 0.979 | 0.969 | 0.917 | 0.969 |
SABL | 0.99 | 0.885 | 0.844 | 0.979 | 0.938 | 0.979 | 0.781 | 0.948 | 0.958 | 0.896 |
Sparse R-CNN | 0.979 | 0.979 | 0.865 | 0.99 | 0.99 | 0.979 | 0.938 | 0.99 | 0.906 | 0.979 |
DETR | 0.979 | 0.771 | 0.615 | 0.604 | 0.896 | 0.708 | 0.885 | 0.635 | 0.625 | 0.906 |
Conditional DETR | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
DDQ | 0.979 | 0.99 | 0.99 | 0.99 | 0.979 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 |
DAB-DETR | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 1.0 | 0.979 | 0.99 | 0.99 |
Deformable DETR | 0.99 | 0.99 | 0.969 | 0.979 | 0.969 | 0.969 | 0.948 | 1.0 | 0.979 | 0.979 |
DINO | 0.99 | 0.948 | 0.979 | 0.99 | 0.969 | 0.99 | 0.99 | 0.979 | 0.979 | 0.99 |
PVT | 0.958 | 0.979 | 0.792 | 0.958 | 0.969 | 0.979 | 0.875 | 0.979 | 0.969 | 0.958 |
PVTv2 | 0.979 | 0.99 | 0.917 | 1.0 | 1.0 | 0.99 | 0.875 | 0.99 | 0.99 | 0.948 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 1.0 | 0.98 | 0.86 | 0.98 | 0.99 | 1.0 | 0.92 | 1.0 | 0.9 | 0.91 |
AutoAssign | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 |
CenterNet | 0.99 | 1.0 | 0.98 | 1.0 | 1.0 | 1.0 | 0.98 | 1.0 | 0.99 | 0.98 |
CentripetalNet | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 |
CornerNet | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DDOD | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DyHead | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
EfficientNet | 1.0 | 1.0 | 0.98 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
FCOS | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 |
FoveaBox | 1.0 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
FreeAnchor | 1.0 | 0.87 | 0.93 | 0.96 | 0.89 | 1.0 | 0.89 | 1.0 | 0.98 | 0.9 |
FSAF | 1.0 | 0.98 | 0.94 | 0.99 | 0.96 | 1.0 | 0.96 | 1.0 | 0.95 | 0.91 |
GFL | 1.0 | 0.95 | 0.93 | 0.99 | 1.0 | 1.0 | 0.95 | 1.0 | 0.86 | 0.9 |
LD | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.98 | 1.0 | 0.93 | 1.0 |
NAS-FPN | 1.0 | 1.0 | 0.98 | 0.99 | 0.98 | 1.0 | 0.98 | 1.0 | 0.99 | 0.97 |
PAA | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
RetinaNet | 1.0 | 0.97 | 0.98 | 0.99 | 1.0 | 1.0 | 0.96 | 1.0 | 0.91 | 0.89 |
RTMDet | 1.0 | 1.0 | 1.0 | 0.97 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 | 1.0 |
TOOD | 1.0 | 1.0 | 0.96 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
VarifocalNet | 0.99 | 0.94 | 0.97 | 0.98 | 0.99 | 1.0 | 0.94 | 1.0 | 0.88 | 0.91 |
YOLOv5 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv6 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv7 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv8 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOX | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.98 | 1.0 | 1.0 | 1.0 | 1.0 |
Faster R-CNN | 1.0 | 0.86 | 0.77 | 0.95 | 0.91 | 1.0 | 0.83 | 1.0 | 0.83 | 0.85 |
Cascade R-CNN | 1.0 | 0.89 | 0.91 | 0.99 | 0.97 | 1.0 | 0.93 | 0.99 | 0.88 | 0.9 |
Cascade RPN | 1.0 | 1.0 | 0.97 | 1.0 | 1.0 | 0.99 | 0.96 | 1.0 | 1.0 | 0.93 |
Double Heads | 1.0 | 1.0 | 0.95 | 1.0 | 0.97 | 1.0 | 0.95 | 1.0 | 0.92 | 0.91 |
FPG | 1.0 | 1.0 | 0.96 | 0.99 | 1.0 | 0.99 | 0.97 | 1.0 | 1.0 | 0.99 |
Grid R-CNN | 0.98 | 0.97 | 0.89 | 1.0 | 0.95 | 0.99 | 0.95 | 1.0 | 0.94 | 0.9 |
Guided Anchoring | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 |
HRNet | 1.0 | 0.98 | 1.0 | 0.99 | 0.94 | 0.99 | 0.84 | 1.0 | 0.98 | 0.95 |
Libra R-CNN | 1.0 | 1.0 | 0.91 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 | 0.96 | 0.96 |
PAFPN | 1.0 | 0.93 | 0.83 | 0.97 | 0.99 | 1.0 | 0.95 | 1.0 | 0.9 | 0.9 |
RepPoints | 1.0 | 1.0 | 0.89 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 0.96 | 0.95 |
Res2Net | 0.98 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 0.93 | 1.0 |
ResNeSt | 0.96 | 0.85 | 0.94 | 0.99 | 1.0 | 0.96 | 0.99 | 0.99 | 0.9 | 0.99 |
SABL | 1.0 | 1.0 | 0.89 | 0.99 | 0.96 | 1.0 | 0.92 | 1.0 | 0.89 | 0.89 |
Sparse R-CNN | 1.0 | 1.0 | 0.88 | 0.99 | 1.0 | 0.98 | 1.0 | 1.0 | 0.92 | 0.95 |
DETR | 1.0 | 0.84 | 0.72 | 0.49 | 0.93 | 0.93 | 0.96 | 0.87 | 0.81 | 0.94 |
Conditional DETR | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DDQ | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
Deformable DETR | 0.99 | 1.0 | 0.99 | 1.0 | 0.99 | 1.0 | 0.99 | 1.0 | 0.98 | 1.0 |
DINO | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.98 | 1.0 |
PVT | 1.0 | 0.98 | 0.7 | 0.98 | 0.96 | 1.0 | 0.92 | 1.0 | 0.96 | 0.85 |
PVTv2 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.82 | 1.0 | 1.0 | 0.97 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.98 | 0.96 | 0.47 | 0.73 | 0.94 | 0.85 | 0.77 | 1.0 | 0.76 | 0.7 |
AutoAssign | 1.0 | 0.86 | 0.74 | 0.92 | 0.91 | 0.97 | 0.71 | 0.93 | 0.87 | 0.98 |
CenterNet | 1.0 | 1.0 | 0.62 | 0.98 | 0.98 | 0.99 | 0.71 | 0.91 | 0.92 | 1.0 |
CentripetalNet | 1.0 | 1.0 | 0.78 | 1.0 | 0.96 | 1.0 | 0.74 | 1.0 | 0.88 | 0.98 |
CornerNet | 0.99 | 0.79 | 0.49 | 0.75 | 0.93 | 0.78 | 0.51 | 0.95 | 0.66 | 0.6 |
DDOD | 0.99 | 0.97 | 0.61 | 0.99 | 0.85 | 0.98 | 1.0 | 1.0 | 0.96 | 0.97 |
DyHead | 1.0 | 0.81 | 0.52 | 0.53 | 0.98 | 0.81 | 0.61 | 0.55 | 0.62 | 0.93 |
EfficientNet | 1.0 | 0.79 | 0.81 | 1.0 | 0.84 | 1.0 | 0.78 | 1.0 | 0.97 | 1.0 |
FCOS | 1.0 | 0.94 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
FoveaBox | 1.0 | 1.0 | 0.49 | 0.96 | 0.87 | 0.99 | 0.67 | 0.87 | 0.96 | 0.92 |
FreeAnchor | 1.0 | 0.82 | 1.0 | 0.98 | 0.9 | 1.0 | 0.99 | 0.98 | 0.88 | 1.0 |
FSAF | 1.0 | 1.0 | 0.64 | 1.0 | 0.82 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 |
GFL | 1.0 | 0.97 | 0.45 | 0.95 | 0.93 | 0.96 | 0.94 | 0.98 | 0.73 | 1.0 |
LD | 1.0 | 0.99 | 0.56 | 1.0 | 1.0 | 0.93 | 0.98 | 0.88 | 0.96 | 1.0 |
NAS-FPN | 1.0 | 0.92 | 0.52 | 0.86 | 1.0 | 0.79 | 0.72 | 0.94 | 0.65 | 0.92 |
PAA | 1.0 | 1.0 | 0.97 | 0.99 | 0.94 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 |
RetinaNet | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
RTMDet | 1.0 | 1.0 | 0.99 | 0.98 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.88 |
TOOD | 0.83 | 0.85 | 0.52 | 0.87 | 0.76 | 0.84 | 0.8 | 0.97 | 0.96 | 0.8 |
VarifocalNet | 1.0 | 0.77 | 0.48 | 0.55 | 0.86 | 0.83 | 0.64 | 0.62 | 0.61 | 0.65 |
YOLOv5 | 1.0 | 1.0 | 0.74 | 1.0 | 1.0 | 1.0 | 0.92 | 1.0 | 1.0 | 1.0 |
YOLOv6 | 1.0 | 1.0 | 0.73 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv7 | 0.98 | 0.76 | 0.78 | 0.72 | 0.91 | 1.0 | 0.78 | 1.0 | 0.75 | 0.79 |
YOLOv8 | 1.0 | 0.92 | 0.64 | 0.82 | 1.0 | 0.82 | 0.64 | 1.0 | 0.83 | 0.97 |
YOLOX | 1.0 | 0.9 | 0.62 | 0.92 | 0.95 | 0.99 | 0.88 | 1.0 | 0.97 | 0.98 |
Faster R-CNN | 0.85 | 0.49 | 0.46 | 0.52 | 0.66 | 0.73 | 0.53 | 0.63 | 0.55 | 0.49 |
Cascade R-CNN | 0.75 | 0.64 | 0.47 | 0.54 | 0.67 | 0.67 | 0.62 | 0.65 | 0.55 | 0.6 |
Cascade RPN | 1.0 | 0.92 | 0.51 | 0.79 | 0.88 | 1.0 | 0.96 | 1.0 | 0.85 | 0.82 |
Double Heads | 0.76 | 0.72 | 0.46 | 0.57 | 0.68 | 0.71 | 0.63 | 0.87 | 0.6 | 0.62 |
FPG | 1.0 | 0.89 | 0.5 | 0.93 | 0.97 | 0.9 | 0.72 | 0.91 | 0.96 | 0.53 |
Grid R-CNN | 0.91 | 0.68 | 0.48 | 0.68 | 0.71 | 0.77 | 0.63 | 0.85 | 0.6 | 0.56 |
Guided Anchoring | 1.0 | 1.0 | 0.87 | 0.99 | 1.0 | 1.0 | 0.8 | 1.0 | 1.0 | 0.9 |
HRNet | 0.96 | 0.78 | 0.55 | 0.59 | 0.73 | 0.83 | 0.62 | 0.81 | 0.62 | 0.62 |
Libra R-CNN | 1.0 | 1.0 | 0.67 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
PAFPN | 0.74 | 0.63 | 0.4 | 0.51 | 0.63 | 0.71 | 0.62 | 0.66 | 0.57 | 0.44 |
RepPoints | 1.0 | 1.0 | 0.63 | 1.0 | 0.94 | 0.77 | 1.0 | 1.0 | 1.0 | 0.9 |
Res2Net | 0.9 | 0.64 | 0.57 | 0.55 | 0.73 | 0.65 | 0.66 | 0.77 | 0.66 | 0.72 |
ResNeSt | 0.97 | 0.61 | 0.5 | 0.73 | 0.72 | 0.96 | 0.61 | 0.93 | 0.61 | 0.58 |
SABL | 0.77 | 0.53 | 0.45 | 0.54 | 0.66 | 0.66 | 0.61 | 0.66 | 0.56 | 0.47 |
Sparse R-CNN | 1.0 | 0.96 | 0.63 | 0.91 | 1.0 | 0.72 | 0.8 | 0.78 | 0.87 | 0.85 |
DETR | 0.77 | 0.56 | 0.39 | 0.45 | 0.71 | 0.64 | 0.51 | 0.55 | 0.55 | 0.57 |
Conditional DETR | 1.0 | 0.88 | 0.67 | 0.84 | 1.0 | 0.8 | 0.86 | 1.0 | 0.84 | 0.95 |
DDQ | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 | 0.98 |
Deformable DETR | 1.0 | 1.0 | 0.94 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DINO | 1.0 | 1.0 | 0.93 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
PVT | 1.0 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.96 |
PVTv2 | 1.0 | 0.93 | 0.64 | 0.94 | 0.87 | 1.0 | 0.76 | 1.0 | 0.76 | 0.6 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.97 | 0.71 | 0.45 | 0.79 | 0.84 | 0.85 | 0.68 | 0.98 | 0.71 | 0.73 |
AutoAssign | 0.98 | 0.9 | 0.74 | 0.92 | 0.94 | 1.0 | 0.83 | 0.98 | 0.91 | 0.99 |
CenterNet | 0.98 | 0.91 | 0.58 | 0.91 | 0.92 | 0.89 | 0.93 | 0.9 | 0.88 | 0.94 |
CentripetalNet | 0.98 | 0.78 | 0.66 | 0.79 | 0.9 | 0.86 | 0.7 | 0.85 | 0.71 | 0.86 |
CornerNet | 0.96 | 0.75 | 0.61 | 0.87 | 0.93 | 0.86 | 0.7 | 0.94 | 0.76 | 0.82 |
DDOD | 0.99 | 0.96 | 0.74 | 0.99 | 0.86 | 1.0 | 0.82 | 1.0 | 0.97 | 0.94 |
DyHead | 1.0 | 0.79 | 0.53 | 0.79 | 0.92 | 0.8 | 0.79 | 0.73 | 0.77 | 0.99 |
EfficientNet | 1.0 | 0.94 | 0.86 | 0.99 | 0.98 | 1.0 | 0.98 | 0.98 | 0.96 | 0.93 |
FCOS | 1.0 | 0.99 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
FoveaBox | 0.94 | 0.84 | 0.5 | 0.76 | 0.87 | 0.81 | 0.72 | 0.8 | 0.76 | 0.91 |
FreeAnchor | 0.97 | 0.73 | 0.79 | 0.88 | 0.81 | 0.98 | 0.74 | 0.85 | 0.84 | 0.88 |
FSAF | 0.93 | 0.75 | 0.49 | 0.8 | 0.72 | 0.81 | 0.67 | 0.89 | 0.74 | 0.84 |
GFL | 1.0 | 0.8 | 0.48 | 0.9 | 0.95 | 0.95 | 0.89 | 0.92 | 0.81 | 0.9 |
LD | 0.98 | 0.86 | 0.52 | 0.9 | 0.92 | 0.87 | 0.75 | 0.94 | 0.85 | 0.93 |
NAS-FPN | 1.0 | 0.92 | 0.66 | 0.96 | 0.98 | 0.97 | 0.73 | 0.82 | 0.75 | 0.88 |
PAA | 0.99 | 0.96 | 0.9 | 0.97 | 0.93 | 0.99 | 0.86 | 0.99 | 1.0 | 0.95 |
RetinaNet | 1.0 | 0.81 | 0.67 | 0.85 | 0.89 | 0.86 | 0.81 | 0.84 | 0.83 | 0.8 |
RTMDet | 1.0 | 1.0 | 0.93 | 0.99 | 1.0 | 0.98 | 0.99 | 1.0 | 1.0 | 0.98 |
TOOD | 0.9 | 0.72 | 0.48 | 0.9 | 0.73 | 0.79 | 0.72 | 0.84 | 0.74 | 0.86 |
VarifocalNet | 0.99 | 0.75 | 0.45 | 0.66 | 0.87 | 0.83 | 0.67 | 0.8 | 0.67 | 0.79 |
YOLOv5 | 1.0 | 1.0 | 0.98 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.97 | 0.99 |
YOLOv6 | 1.0 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 |
YOLOv7 | 1.0 | 0.94 | 0.97 | 0.94 | 0.98 | 1.0 | 0.96 | 1.0 | 0.98 | 0.99 |
YOLOv8 | 1.0 | 0.95 | 0.87 | 0.93 | 1.0 | 0.98 | 0.89 | 0.99 | 0.84 | 0.99 |
YOLOX | 1.0 | 0.97 | 0.77 | 0.97 | 0.99 | 1.0 | 0.98 | 0.92 | 0.94 | 0.98 |
Faster R-CNN | 0.85 | 0.45 | 0.35 | 0.47 | 0.54 | 0.59 | 0.49 | 0.61 | 0.44 | 0.49 |
Cascade R-CNN | 0.82 | 0.51 | 0.4 | 0.52 | 0.58 | 0.64 | 0.53 | 0.62 | 0.48 | 0.51 |
Cascade RPN | 1.0 | 0.91 | 0.62 | 0.93 | 0.96 | 0.99 | 0.96 | 1.0 | 0.86 | 0.96 |
Double Heads | 0.83 | 0.61 | 0.44 | 0.58 | 0.58 | 0.72 | 0.54 | 0.8 | 0.57 | 0.57 |
FPG | 1.0 | 0.78 | 0.5 | 0.93 | 0.82 | 0.99 | 0.7 | 0.94 | 0.9 | 0.71 |
Grid R-CNN | 0.89 | 0.51 | 0.37 | 0.6 | 0.58 | 0.73 | 0.55 | 0.76 | 0.47 | 0.55 |
Guided Anchoring | 1.0 | 0.97 | 0.98 | 0.99 | 0.99 | 1.0 | 0.89 | 1.0 | 0.97 | 0.97 |
HRNet | 0.89 | 0.6 | 0.52 | 0.59 | 0.54 | 0.73 | 0.51 | 0.84 | 0.56 | 0.6 |
Libra R-CNN | 0.94 | 0.76 | 0.49 | 0.8 | 0.9 | 0.74 | 0.85 | 0.81 | 0.72 | 0.93 |
PAFPN | 0.84 | 0.55 | 0.37 | 0.52 | 0.59 | 0.65 | 0.58 | 0.71 | 0.5 | 0.54 |
RepPoints | 0.97 | 0.89 | 0.52 | 0.84 | 0.83 | 0.84 | 0.8 | 0.95 | 0.94 | 0.84 |
Res2Net | 0.88 | 0.59 | 0.52 | 0.66 | 0.77 | 0.83 | 0.54 | 0.79 | 0.55 | 0.73 |
ResNeSt | 0.91 | 0.63 | 0.43 | 0.8 | 0.78 | 0.91 | 0.68 | 0.92 | 0.64 | 0.62 |
SABL | 0.82 | 0.49 | 0.4 | 0.52 | 0.56 | 0.57 | 0.53 | 0.71 | 0.46 | 0.49 |
Sparse R-CNN | 0.99 | 0.89 | 0.55 | 0.79 | 0.95 | 0.81 | 0.83 | 0.9 | 0.82 | 0.86 |
DETR | 0.71 | 0.44 | 0.3 | 0.32 | 0.57 | 0.52 | 0.63 | 0.43 | 0.38 | 0.5 |
Conditional DETR | 1.0 | 0.9 | 0.71 | 0.94 | 1.0 | 0.92 | 0.8 | 0.98 | 0.9 | 0.94 |
DDQ | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 0.94 | 1.0 | 0.99 | 0.99 | 1.0 | 0.98 | 0.96 | 1.0 | 0.99 |
Deformable DETR | 1.0 | 0.9 | 0.74 | 0.86 | 0.87 | 0.93 | 0.95 | 0.96 | 0.86 | 0.91 |
DINO | 1.0 | 0.92 | 0.95 | 0.99 | 0.96 | 1.0 | 1.0 | 1.0 | 0.92 | 1.0 |
PVT | 0.98 | 0.95 | 0.78 | 0.87 | 0.96 | 0.97 | 0.91 | 0.94 | 0.87 | 0.8 |
PVTv2 | 1.0 | 0.99 | 0.77 | 0.97 | 1.0 | 1.0 | 0.78 | 0.99 | 0.98 | 0.8 |
Clean | AdvCam | RP2 | ShapeShifter | |
ATSS | 0.929 | 0.93 | 0.892 | 0.919 |
AutoAssign | 0.921 | 0.943 | 0.896 | 0.915 |
CenterNet | 0.903 | 0.91 | 0.875 | 0.914 |
CentripetalNet | 0.951 | 0.946 | 0.924 | 0.951 |
CornerNet | 0.942 | 0.951 | 0.929 | 0.947 |
DDOD | 0.916 | 0.926 | 0.9 | 0.915 |
DyHead | 0.927 | 0.933 | 0.866 | 0.921 |
EfficientNet | 0.921 | 0.913 | 0.887 | 0.919 |
FCOS | 0.939 | 0.932 | 0.913 | 0.929 |
FoveaBox | 0.915 | 0.917 | 0.871 | 0.913 |
FreeAnchor | 0.921 | 0.919 | 0.877 | 0.912 |
FSAF | 0.901 | 0.907 | 0.864 | 0.899 |
GFL | 0.927 | 0.942 | 0.887 | 0.908 |
LD | 0.933 | 0.931 | 0.88 | 0.92 |
NAS-FPN | 0.93 | 0.942 | 0.883 | 0.925 |
PAA | 0.928 | 0.921 | 0.886 | 0.91 |
RetinaNet | 0.921 | 0.912 | 0.863 | 0.899 |
RTMDet | 0.929 | 0.942 | 0.862 | 0.927 |
TOOD | 0.922 | 0.93 | 0.895 | 0.921 |
VarifocalNet | 0.929 | 0.932 | 0.902 | 0.921 |
YOLOv5 | 0.941 | 0.943 | 0.882 | 0.945 |
YOLOv6 | 0.94 | 0.954 | 0.901 | 0.936 |
YOLOv7 | 0.945 | 0.942 | 0.89 | 0.942 |
YOLOv8 | 0.942 | 0.943 | 0.868 | 0.944 |
YOLOX | 0.923 | 0.922 | 0.866 | 0.906 |
Faster R-CNN | 0.891 | 0.897 | 0.863 | 0.861 |
Cascade R-CNN | 0.929 | 0.924 | 0.891 | 0.895 |
Cascade RPN | 0.927 | 0.93 | 0.887 | 0.901 |
Double Heads | 0.887 | 0.895 | 0.847 | 0.88 |
FPG | 0.921 | 0.935 | 0.859 | 0.897 |
Grid R-CNN | 0.913 | 0.911 | 0.865 | 0.91 |
Guided Anchoring | 0.928 | 0.921 | 0.899 | 0.925 |
HRNet | 0.923 | 0.915 | 0.91 | 0.904 |
Libra R-CNN | 0.921 | 0.922 | 0.885 | 0.905 |
PAFPN | 0.901 | 0.89 | 0.85 | 0.882 |
RepPoints | 0.915 | 0.908 | 0.865 | 0.887 |
Res2Net | 0.91 | 0.897 | 0.861 | 0.899 |
ResNeSt | 0.929 | 0.901 | 0.872 | 0.889 |
SABL | 0.92 | 0.913 | 0.88 | 0.898 |
Sparse R-CNN | 0.931 | 0.925 | 0.9 | 0.927 |
DETR | 0.908 | 0.904 | 0.878 | 0.933 |
Conditional DETR | 0.93 | 0.931 | 0.907 | 0.924 |
DDQ | 0.949 | 0.95 | 0.901 | 0.932 |
DAB-DETR | 0.931 | 0.94 | 0.902 | 0.919 |
Deformable DETR | 0.943 | 0.955 | 0.893 | 0.933 |
DINO | 0.94 | 0.944 | 0.885 | 0.936 |
PVT | 0.906 | 0.886 | 0.868 | 0.861 |
PVTv2 | 0.915 | 0.899 | 0.877 | 0.905 |
B.5 Additional ablation experiments
Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 12, 13, 14, 15, 16, and 17. In addition, the visualized evaluation results are shown in Fig. 16, 17, and 18.
Appendix C Reproducibility
The reproducibility of all evaluation results in our benchmark is ensured through the provision of the necessary code in the associated Github Repository (https://github.com/JiaweiLian/Benchmarking_Physical_Attack). Furthermore, all datasets and model architectures evaluated are accessible to the public.
Appendix D License
This work is under Apache-2.0 license. More detailed information can be found in the LICENSE file in the GitHub repository: https://github.com/JiaweiLian/Benchmarking_Physical_Attack